### v1

# Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1

Abstract
Event cameras offer unparalleled advantages for real-time perception in dynamic environments, thanks to their microsecond-level temporal resolution and asynchronous operation. Existing event-based object detection methods, however, are limited by fixed-frequency paradigms and fail to fully exploit the high-temporal resolution and adaptability of event cameras. To address these limitations, we propose FlexEvent, a novel event camera object detection framework that enables detection at arbitrary frequencies. Our approach consists of two key components: FlexFuser, an adaptive event-frame fusion module that integrates high-frequency event data with rich semantic information from RGB frames, and FAL, a frequency-adaptive learning mechanism that generates frequency-adjusted labels to enhance model generalization across varying operational frequencies. This combination allows our method to detect objects with high accuracy in both fast-moving and static scenarios, while adapting to dynamic environments. Extensive experiments on large-scale event camera datasets demonstrate that our approach surpasses state-of-the-art methods, achieving significant improvements in both standard and high-frequency settings. Notably, our method maintains robust performance when scaling from $20$ Hz to $90$ Hz and delivers accurate detection up to $180$ Hz, proving its effectiveness in extreme conditions. Our framework sets a new benchmark for event-based object detection and paves the way for more adaptable, real-time vision systems.
Figure 1: Event camera detection at varying frequencies. The performance of the classic RVT detector (Gehrig & Scaramuzza, 2023) drops significantly at higher event operational frequencies. Motivated by this, we propose FlexEvent, a robust and flexible detector that maintain high accuracy across a wide range of frequencies, ensuring strong adaptability in dynamic sensing environments.
1 Introduction
Event cameras have garnered significant attention for their ability to capture dynamic scenes with microsecond-level temporal resolution (Gallego et al., 2022). Unlike conventional RGB cameras, which capture entire frames at fixed intervals, event cameras operate asynchronously, responding to changes in pixel intensity at each location (Zou et al., 2022). This low-latency operation reduces motion blur and enables highly energy-efficient sensing, making event cameras ideal for real-time applications such as autonomous driving, robotics, and surveillance (Steffen et al., 2019; Kong et al., 2024b; Chen et al., 2023).
Despite their potential, existing event camera object detection methods often fail to fully leverage the high-frequency temporal information captured by event cameras (Cordone et al., 2022; Gehrig & Scaramuzza, 2022; Jeziorek et al., 2023). Most approaches align event data with the lower frequency of RGB cameras by adopting a fixed time interval between event streams and frame-based annotations (Perot et al., 2020; Gehrig et al., 2021b). While this strategy simplifies data processing, it inevitably overlooks the rich temporal details embedded in high-frequency event streams. Given that human annotations are often synchronized with slower frame rates, current detection models miss valuable information from high-frequency event data, resulting in suboptimal performance when rapid object detection is required in dynamic environments (Messikommer et al., 2020; Schaefer et al., 2022).
To address these limitations, we introduce FlexEvent, a novel event camera object detection framework designed to tackle the challenging problem of object detection at varying operational frequencies. Our approach addresses the need for high-frequency detection in fast-changing environments, while adapting to different operational frequencies. We propose two key innovations: (1) FlexFuser, an adaptive event-frame fusion module, and (2) FAL, a frequency-adaptive learning mechanism.
Flexible Event-Frame Fusion. The first component, FlexFuser, addresses the limitations of event data, which often lacks semantic and texture-rich information, especially at higher frequencies (Zhou et al., 2023), by integrating the rich spatial and semantic information from RGB frames with the high-temporal resolution of event streams. It enables high detection accuracy even in fast-moving environments. Furthermore, training on high-frequency event data is computationally expensive and impractical due to the significant human effort required to label such data. FlexFuser mitigates this by sampling event data at varying frequencies, aligning them with the normal frame rate during training, thus maintaining efficiency while preserving the high-frequency benefits at inference time.
Frequency-Adaptive Learning. The second component, FAL, enhances the generalization capability of event camera detectors across varying operational frequencies, by generating frequency-adjusted labels for the unlabeled high-frequency data. These labels allow the model to learn from high-frequency event streams without manual annotations, and iterative refinement through self-training ensures that the model remains robust across different motion dynamics and frequency settings. Together, these two components allow for accurate real-time detection in rapid scene changes and adapt to a wide range of operational frequencies, by leveraging the temporal richness of event data and the semantic detail of RGB frames.
Our extensive experiments validate the effectiveness of FlexEvent on multiple large-scale event camera datasets. Our approach consistently outperforms recent detectors across both standard and high-frequency settings. In particular, we achieve mAP gains of $\mathbf{15.5\%}$, $\mathbf{9.4\%}$, and $\mathbf{10.3\%}$ over the previous best-performing detectors on the DSEC-Det (Gehrig & Scaramuzza, 2024), DSEC-Detection (Tomy et al., 2022), and DSEC-MOD (Zhou et al., 2023) datasets, respectively. Our model also maintains $\mathbf{96.2\%}$ of its performance when the operational frequency shifts from $\mathbf{20}$ Hz to $\mathbf{90}$ Hz, and delivers accurate detection at frequencies as high as $\mathbf{180}$ Hz, proving its robustness under extreme conditions.
[…]
2 Related Work
Event Camera Object Detection. Event-based detection methods can be broadly split into two approaches: GNNs/SNNs and dense feed-forward models. GNNs build dynamic spatio-temporal graphs by subsampling events (Gehrig & Scaramuzza, 2022; Sun & Ji, 2023; Messikommer et al., 2020; Schaefer et al., 2022), but they face challenges in propagating information over large spatio-temporal regions, especially for slow-moving objects. SNNs offer efficient sparse information transmission but are often hindered by their non-differentiable nature, complicating optimization processes (Cuadrado et al., 2023; Cordone et al., 2022; Zhang et al., 2022). Dense, feed-forward models represent the second approach. Initial methods using fixed temporal windows (Chen, 2018; Iacono et al., 2018; Jiang et al., 2019) struggle with slow-moving or stationary objects due to their limited capability to capture long-term temporal data. Subsequent advancements incorporate RNNs and transformers to enhance temporal modeling capabilities (Perot et al., 2020; Zubić et al., 2023; Li et al., 2022; Gehrig & Scaramuzza, 2023; Peng et al., 2024), but these models often still lack semantic richness and face difficulties in adapting to variable frequencies and highly dynamic scenarios.
Event-Frame Multimodal Learning. To overcome the limited texture in event streams, multimodal fusion techniques combining event-based and frame-based data have gained traction across tasks, such as deblurring (Sun et al., 2022a; Zhang et al., 2020), depth estimation (Gehrig et al., 2021a; Uddin et al., 2022), and tracking (Zhao et al., 2022; Gehrig et al., 2020). Earlier object detection approaches fuse event and image data during post-processing (Li et al., 2019; Chen et al., 2019), but they lack meaningful feature-level interaction. Recent works focus on deeper feature fusion (Tomy et al., 2022; Cao et al., 2022; 2021), with advanced methods introducing pixel-level spatial attention or temporal transformers for asynchronous processing (Zhou et al., 2023; Li et al., 2023; Gehrig & Scaramuzza, 2024; Cao et al., 2024). Some approaches (Li et al., 2023; Gehrig et al., 2021a) explore combining events and frames through asynchronous multi-modal fusion, enabling inference at varying frequencies. However, they do not focus on high-frequency event data and fully leverage the temporally rich nature of event streams. Furthermore, these methods still face challenges in fully exploiting complementary strengths and addressing feature imbalance in event-frame detection. Unlike previous methods, FlexEvent employs a more comprehensive fusion framework that effectively combines high-temporal resolution event data with rich semantics from RGB frames, enabling robust object detection across varying frequencies while addressing feature imbalance.
Label-Efficient Learning in Event Data. Due to limited annotated datasets, label-efficient learning has become an important area for event-based vision. Several studies attempt to reconstruct images from event data (Rebecq et al., 2019; 2021; Stoffregen et al., 2020) or leverage knowledge distillation from pre-trained frame-based models (Wang et al., 2021; Sun et al., 2022b; Yang et al., 2023; Kong et al., 2024a). Other approaches utilize pre-trained models or self-supervised losses (Klenk et al., 2022; Wu et al., 2023; Zhu et al., 2019). LEOD (Wu et al., 2024) pioneers object detection with limited labels but does not address high-frequency generalization. A recent state-space model (Zubić et al., 2024) adapts to varying frequencies without retraining but struggles to detect static objects at high frequencies due to its reliance solely on event data. In contrast, FlexEvent is specifically designed to adapt to varying event frequencies, ensuring consistent performance even in scenarios with limited labels, and effectively detecting both stationary and fast-moving objects.
3 FlexEvent: A Flexible Event Object Detector
In this section, we elaborate on the technical details of our FlexEvent framework. We start with the foundational concepts of event data and their representation in Sec. 3.1. We then introduce the FlexFuser module in Sec. 3.2, which adaptively fuses event and frame data to enhance detection across varying frequencies. Finally, we detail the frequency-adaptive learning (FAL) mechanism in Sec. 3.3, which enables our model to generalize effectively across diverse temporal conditions using self-training and adaptive label generation. The overall framework is illustrated in Fig.2.
Figure 2: Framework Overview. The proposed FlexEvent consists of two branches: Event and Frame. The event branch captures high-temporal resolution data, while the frame branch leverages the rich semantic information from frames (cf. Sec. 3.1). These branches are fused dynamically through FlexFuser, allowing adaptive integration of event and frame data (cf. Sec. 3.2). Additionally, the frequency-adaptive learning (FAL) mechanism ensures robust detection performance across varying operational frequencies (cf. Sec. 3.3). Together, these components enable the model to handle diverse motion dynamics and maintain high detection accuracy in varying frequency scenarios.
3.1 Background & Preliminaries
Event Processing. Event cameras are bio-inspired vision sensors that capture changes in log intensity per pixel asynchronously, rather than capturing entire frames at fixed intervals. Formally, let $I(x,y,t)$ denotes the log intensity at pixel coordinates $(x,y)$ and time $t$. An event $e$ is generated at $(x,y,t)$ whenever the change in log intensity $\Delta I$ exceeds a certain threshold $C$. Such a process can be modeled as:
$\Delta I(x,y,t)=I(x,y,t)-I(x,y,t-\Delta t)\geq C~{}.$ (1)
Each event $e$ is a tuple $(x,y,t,p)$, where $(x,y)$ are the pixel coordinates, $t$ is the timestamp, and $p\in\{-1,1\}$ denotes the polarity of the event which indicates the direction of the intensity change.
To leverage event data with convolutional neural network layers, we preprocess events into a 4D tensor $E$ with dimensions representing the polarity, temporal discretization $T$, and spatial dimensions $(H,W)$. This representation involves mapping a set of events $\mathcal{E}$ within time interval $[t_{a},t_{b})$ into:
$E(p,\tau,x,y)=\sum_{e_{k}\in\mathcal{E}}\delta(p-p_{k})\delta(x-x_{k},y-y_{k})\delta(\tau-\tau_{k}),$ (2)
where $\tau_{k}=\left\lfloor\frac{t_{k}-t_{a}}{t_{b}-t_{a}}\cdot T\right\rfloor$. This tensor captures event activity in $T$ discrete time slices, yielding a compact representation suitable for 2D convolutions by flattening the polarity and temporal dimensions.
Problem Formulation. Given two consecutive frames $F_{a}$ and $F_{b}$ captured at timestamps $T_{a}$ and $T_{b}$, our objective is to leverage the event stream over the interval $[T_{a},T_{b}]$ to detect objects at the end timestamp $T_{b}$. Existing event-based object detection methods often use fixed time intervals $\Delta T$, limiting adaptability to dynamic environments (Perot et al., 2020). Additionally, integrating semantic information from RGB frames remains challenging, affecting performance in complex scenarios (De Tournemire et al., 2020). To address this, we synchronize the event data with frames and explore varying training frequencies, leveraging the temporal richness of event cameras to improve detection accuracy.
3.2 FlexFuser: An Adaptive Event-Frame Fusion Module
In dynamic environments, object detection systems must adapt to varying motion frequencies (Sun et al., 2022a). While event cameras excel at capturing rapid changes in pixel intensity, they often lack the rich spatial and semantic information provided by frames. To address this limitation and fully leverage the complementary strengths of both modalities, we introduce FlexFuser, an adaptive fusion module designed to dynamically combine event data at different frequencies with frame data.
Dynamic Event Aggregation. Given a dataset $\mathcal{D}$, consisting of sequences of calibrated event camera data and frame data with a resolution of $H\times W$, along with corresponding bounding box annotations $y$ collected at frequency $a$, we begin by selecting a batch of frame data $\mathbf{F}$ paired with event data $\mathbf{E}^{a}$, both captured at frequency $a$. To aggregate event data from a higher frequency $b$ (where $b>a$), we divide the time interval $\Delta T^{a}$ corresponding to $\mathbf{E}^{a}$ into $b/a$ smaller sub-intervals. From each sub-interval, we obtain a high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, as defined in Eq. 2.
From this set, we randomly sample one event data point [^1] For simplicity, we use $\mathbf{E}^{b}$ to represent a sample from the set of high-frequency event data $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, rather than explicitly referencing each individual sample from the event set. The same applies to other frequencies. $\mathbf{E}^{b}$. By doing so, the sampled high-frequency event data $\mathbf{E}^{b}$ is temporally aligned with the frame data $\mathbf{F}$ and the base frequency event data $\mathbf{E}^{a}$. This synchronization of event streams at different frequencies ensures consistent and reliable processing for subsequent stages.
Figure 3: Illustration of the FlexFuser module. We show a general example of event and frame under frequency $a$ at stage $i$.
Feature Extraction. Let $\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}(\cdot)$ and $\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}(\cdot)$ represent the event- and frame-based networks, respectively, where the former employs the RVT (Gehrig & Scaramuzza, 2023) for extracting features from event data, and the latter uses ResNet-50 (He et al., 2016) for feature extraction from frames. Both networks are structured into four stages, as shown in Fig. 2.
At each scale $i$, we extract features ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a},^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}$ from the event data and ${}^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}$ from the frame data:
${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}=\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{(i)}(\mathbf{E}^{a}),\quad^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}=\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{(i)}(\mathbf{E}^{b}),\quad^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}=\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}^{(i)}(\mathbf{F}),$ (3)
where ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}$ and ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}\in\mathbb{R}^{B\times C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}\times H_{i}\times W_{i}}$, ${}^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\in\mathbb{R}^{B\times C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\times H_{i}\times W_{i}}$. Here, $i$ denotes the scale, $B$ is the batch size, and $C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}$ and $C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}$ are the dimensions of the feature maps extracted from the event and frame data, respectively.
Event-Frame Adaptive Fuser. To effectively fuse the event and frame data, we employ an adaptive fuser that is consistent across different event data frequencies. At each scale $i$, taking the low-frequency event features ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}$ as an example, we concatenate the feature maps from both the event and frame branches as follows:
${}^{(i)}\mathbf{h}_{\text{shared}}^{a}=\left[{}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a},\ ^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right]\in\mathbb{R}^{B\times\left(C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right)\times H_{i}\times W_{i}}~{}.$ (4)
Inspired by previous works (Zhou et al., 2023) and (Zhong et al., 2024), our goal is to dynamically fuse these two modalities in a flexible manner. The proposed FlexFuser module computes adaptive soft weights that regulate the contribution of each branch (event and frame) based on the current input conditions. As shown in Fig. 3, these adaptive soft weights are computed using a gating function, which incorporates learned noise to introduce perturbations for improved adaptability. The process is:
$\left[{}^{(i)}\bm{\alpha},\ ^{(i)}\bm{\beta}\right]=\text{Softmax}\left((^{(i)}\mathbf{h}_{\text{shared}}^{a}\cdot^{(i)}\mathbf{W})+^{(i)}\sigma\cdot\epsilon\right),$ (5)
where ${}^{(i)}\mathbf{W}\in\mathbb{R}^{\left(C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right)\times 2}$ is a trainable weight matrix, ${}^{(i)}\bm{\alpha}$ and ${}^{(i)}\bm{\beta}$ are the adaptive soft weights for the event and frame branches, respectively. Here, ${}^{(i)}\sigma$ is a learned standard deviation that controls the magnitude of the noise perturbation, and $\epsilon\sim\mathcal{N}(0,1)$ represents a Gaussian noise term.
The fused feature map at each scale $i$ is then obtained by applying the adaptive soft weights to the event and frame features:
${}^{(i)}\mathbf{h}_{\text{fuse}}^{a}=^{(i)}\bm{\alpha}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}+^{(i)}\bm{\beta}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}},$ (6)
where $\odot$ denotes element-wise multiplication. This fusion process dynamically balances the contribution of each modality based on the input data, allowing for more robust and adaptive feature representation across varying conditions.
Then, at each scale $i$, the final feature map combining event data at different frequencies and the frame data is obtained by adding the fused features from different frequencies. Specifically, we combine the fused feature maps as ${}^{(i)}\mathbf{h}_{\text{fuse}}=^{(i)}\mathbf{h}_{\text{fuse}}^{a}+^{(i)}\mathbf{h}_{\text{fuse}}^{b}$. After obtaining the fused feature maps across all scales, the multi-scale features are concatenated and fed into the detection head to produce the predicted bounding box $\mathbf{\hat{y}}$.
Optimization & Regularization. In addition to the standard detection loss $\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}})$, we introduce a regularization term to ensure balanced utilization of both the event and frame branches. This term penalizes large variations in the soft weights, encouraging a more uniform contribution from both modalities and preventing overfitting to a single branch:
$\mathcal{L}_{\text{fuser}}=\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}})+\lambda\left(\frac{\text{Var}\left(\bm{\alpha}\right)}{\left(\mathbb{E}\left[\bm{\alpha}\right]\right)^{2}}+\frac{\text{Var}\left(\bm{\beta}\right)}{\left(\mathbb{E}\left[\bm{\beta}\right]\right)^{2}}\right),$ (7)
where $\lambda$ is a weighting factor.
3.3 FAL: Frequency-Adaptive Learning Mechanism
FlexFuser aggregates information from different frequencies using labeled low-frequency data. To tune the model adaptively to handle diverse frequencies by leveraging both labeled low-frequency data and unlabeled high-frequency data, we design a flexible frequency-adaptive learning (FAL) mechanism. FAL incorporates multi-frequency information into the training process through iterative self-training. This approach enhances the model’s ability to generalize across varying frequencies, making it more robust in different scenarios. The key steps of the FAL mechanism are as follows:
Pre-Training with Low-Frequency Labels. Rather than training solely at the data collection frequency $a$, we enhance the model’s capability by training it at a higher frequency $b$. To efficiently leverage the available labels, we select only the final event from the high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, which corresponds to the labeled timestamp. This approach allows the model to capture valuable high-frequency temporal information while still utilizing low-frequency labels, improving its temporal understanding and robustness. The training objective is to minimize the detection loss over the labeled data as follows:
$\mathcal{L}_{\text{GT}}=\sum_{(\mathbf{F},\mathbf{E}^{b}_{b/a},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}}).$ (8)
Labels at Higher Frequencies. For the unlabeled data in $\mathcal{D}$ captured at frequency $b$, the pre-trained model generates high-frequency labels $\mathbf{\hat{y}}$ by performing inference on the entire high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$. These generated labels $\mathbf{\hat{y}}$ serve as pseudo-labels for guiding further training at higher frequencies, improving the model’s ability to generalize across different temporal conditions.
Enhanced Temporal Refinement. To refine the high-frequency labels, we introduce a multi-step temporal refinement approach. First, we adopt bidirectional event augmentation by processing both forward and reversed event streams to detect objects with varying movements and orientations, thereby boosting recall. After generating the bidirectional labels, we apply Non-Maximum Suppression (NMS) to remove overlapping bounding boxes, followed by a low confidence threshold $\mathbf{\tau}$ to retain potential objects and further improve recall. Next, leveraging a tracking-by-detection framework, we link detection boxes across frames using pairwise IoU matching with a threshold $\mathbf{\tau}^{\text{IoU}}$. Short-lived tracks, with lengths below $\mathbf{L}^{\text{track}}$, are pruned to ensure temporal consistency. This approach ensures that the refined high-frequency labels $\tilde{\mathbf{y}}$ are accurate, temporally consistent, and reliable, ultimately improving detection quality in high-frequency data even in the absence of ground truth labels.
Self-Training Iteration. The model is iteratively trained using the refined high-frequency labels $\tilde{\mathbf{y}}$ on high-frequency data where no ground truth labels are available. The total loss function combines the base training loss and the pseudo-label loss as:
$\mathcal{L}_{\text{FAL}}=\mathcal{L}_{\text{GT}}+\beta\sum_{(\mathbf{F},\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a-1},\mathbf{\tilde{y}})\in\mathcal{D}}\mathcal{L}_{\text{det}}\left(\mathbf{\tilde{y}},\mathbf{\hat{y}}\right)~{},$ (9)
where $\beta$ balances the contribution of the high-frequency label loss. The complete FlexEvent framework combines FlexFuser and FAL, allowing the model to dynamically fuse event and frame data while adapting to varying frequencies. As demonstrated in the following sections, this combination provides a robust detection framework capable of maintaining high accuracy in dynamic environments.
4 Experiments
Table 1: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). Both event-only and event-frame fusion methods are compared. The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
Modality Method Venue Reference mAP AP50 AP75 APS APM APL
E RVT CVPR’23 (Gehrig & Scaramuzza, 2023) $38.4\%$ $58.7\%$ $41.3\%$ $29.5\%$ $50.3\%$ $81.7\%$
SAST CVPR’24 (Peng et al., 2024) $38.1\%$ $60.1\%$ $40.0\%$ $29.8\%$ $48.9\%$ $79.7\%$
SSM CVPR’24 (Zubić et al., 2024) $38.0\%$ $55.2\%$ $40.6\%$ $28.8\%$ $52.2\%$ $77.8\%$
LEOD CVPR’24 (Wu et al., 2024) $41.1\%$ $65.2\%$ $43.6\%$ $35.1\%$ $47.3\%$ $73.3\%$
E + F DAGr-18 Nature’24 (Gehrig & Scaramuzza, 2024) $37.6\%$
DAGr-34 Nature’24 (Gehrig & Scaramuzza, 2024) $39.0\%$
DAGr-50 Nature’24 (Gehrig & Scaramuzza, 2024) $41.9\%$ $66.0\%$ $44.3\%$ $36.3\%$ $56.2\%$ $77.8\%$
FlexEvent Ours - $\mathbf{57.4\%}$ $\mathbf{78.2\%}$ $\mathbf{66.6\%}$ $\mathbf{51.7\%}$ $\mathbf{64.9\%}$ $\mathbf{83.7\%}$
4.1 Experimental Settings
Datasets. We conduct experiments based on three large-scale event camera datasets: 1DSEC-Det (Gehrig & Scaramuzza, 2024), 2DSEC-Detection (Tomy et al., 2022), and 3DSEC-MOD (Zhou et al., 2023). These datasets comprise $78,344$ frames across $60$ sequences, $52,727$ frames over $41$ sequences, and $13,314$ frames within $16$ sequences, respectively, making them suitable for evaluating event-based object detection methods. We prioritize DSEC-Det as the primary benchmark for reporting results as it is the largest, most recent, and most comprehensive event-frame perception dataset. For more details, please refer to the Appendix.
Baselines. To evaluate our method, we compare it against both event-only and event-frame fusion state-of-the-art methods. For event-only detectors, we include RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), LEOD (Wu et al., 2024), and SSM (Zubić et al., 2024), retraining them on the DSEC-Det dataset following their respective protocols. For event-frame fusion methods, we compare with DAGr (Gehrig & Scaramuzza, 2024) on DSEC-Det using results from the original paper and retrain our method on DSEC-Detection and DSEC-MOD using standard settings to compare with CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023). For other methods on DSEC-Detection and DSEC-MOD, we reference results reported in the CAFR and RENet papers. For more details, please refer to the Appendix.
Implementation Details. We train our model using the YOLOX framework (Zheng et al., 2021), optimizing with the standard detection loss, which consists of IoU loss, classification loss, and regression loss, averaged across both the batch and sequence length for stable training. Additionally, we introduce an extra regularization term to balance the utilization among the different modalities. The model is trained for $100,000$ iterations with a batch size of $8$ and a sequence length of $11$, using a learning rate of $1\times 10^{-4}$. All experiments are conducted on two NVIDIA RTX A5000 GPUs with 24GB memory, with the entire training process completed in approximately one day. Due to space limits, more details are included in the Appendix.
Evaluation Metrics. We evaluate object detectors using the mean Average Precision (mAP) as the primary metric, along with AP50, AP75, APS, APM, and APL from the COCO evaluation protocol (Lin et al., 2014). These metrics provide a comprehensive assessment of detection performance across different IoU thresholds and object sizes. Please refer to the Appendix for more details.
4.2 Comparisons to State-of-the-Art Detectors
Comparison to Event-Only Models. We compare FlexEvent with state-of-the-art event-only detectors, including RVT (Gehrig & Scaramuzza, 2023), SSM (Zubić et al., 2024), SAST (Peng et al., 2024), and LEOD (Wu et al., 2024), as shown in Tab. 1. We significantly outperform these methods across all metrics, with the performance gap becoming even more pronounced at higher frequencies. Event-only methods struggle to maintain detection accuracy at higher frequencies due to their inability to fully capture object semantics. In contrast, we overcome these limitations through the FlexFuser module, which integrates RGB data to compensate for the lack of semantic richness in the event stream. By fusing both event and frame data, we excel in complex and dynamic environments, achieving superior detection accuracy where event-only methods fall short.
Table 2: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Detection (Tomy et al., 2022). Both event-only and event-frame fusion methods are compared. The reported results are the mAP scores of 1Car, 2Pedestrian (Ped), and 3Large-Vehicle (L-Veh) classes. The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
Modality Method Venue Reference Type Car Ped L-Veh Average
E CAFR ECCV’24 (Cao et al., 2024) Event - - - $12.0\%$
E + F SENet CVPR’18 (Hu et al., 2018) Attention $38.4\%$ $14.9\%$ $26.0\%$ $26.2\%$
CBAM ECCV’18 (Woo et al., 2018) $37.7\%$ $13.5\%$ $27.0\%$ $26.1\%$
ECA-Net CVPR’20 (Wang et al., 2020) $36.7\%$ $12.8\%$ $27.5\%$ $25.7\%$
SAGate ECCV’20 (Chen et al., 2020) RGB + Depth $32.5\%$ $10.4\%$ $16.0\%$ $19.6\%$
DCF CVPR’21 (Ji et al., 2021) $36.3\%$ $12.7\%$ $28.0\%$ $25.7\%$
SPNet ICCV’21 (Zhou et al., 2021) $39.2\%$ $17.8\%$ $26.2\%$ $27.7\%$
RAMNet RA-L’21 (Gehrig et al., 2021a) $24.4\%$ $10.8\%$ $17.6\%$ $17.6\%$
FAGC Sensors’21 (Cao et al., 2021) $39.8\%$ $14.4\%$ $33.6\%$ $29.3\%$
FPN-Fusion ICRA’22 (Tomy et al., 2022) RGB + Event $37.5\%$ $10.9\%$ $24.9\%$ $24.4\%$
EFNet ECCV’22 (Sun et al., 2022a) $41.1\%$ $15.8\%$ $32.6\%$ $30.0\%$
DRFuser EAAI’23 (Munir et al., 2023) $38.6\%$ $15.1\%$ $30.6\%$ $28.1\%$
CMX TITS’23 (Zhang et al., 2023) $41.6\%$ $16.4\%$ $29.4\%$ $29.1\%$
RENet ICRA’23 (Zhou et al., 2023) $40.5\%$ $17.2\%$ $30.6\%$ $29.4\%$
CAFR ECCV’24 (Cao et al., 2024) $49.9\%$ $25.8\%$ $38.2\%$ $38.0\%$
FlexEvent Ours - $\mathbf{59.3\%}$ $\mathbf{37.4\%}$ $\mathbf{45.5\%}$ $\mathbf{47.4\%}$
Figure 4: Comparisons of event-based detectors under different event frequencies on DSEC-Det.
Table 3: Comparisons of fusion-based event detectors on the test set of DSEC-MOD (Zhou et al., 2023). The best and 2nd best scores from each metric are highlighted in bold and underlined.
Method Venue Reference mAP
SENet CVPR’18 (Hu et al., 2018) $29.28\%$
CBAM ECCV’18 (Woo et al., 2018) $36.22\%$
ECA-Net CVPR’20 (Wang et al., 2020) $34.49\%$
SAGate ECCV’20 (Chen et al., 2020) $33.62\%$
DCF CVPR’21 (Ji et al., 2021) $32.20\%$
SPNet ICCV’21 (Zhou et al., 2021) $32.70\%$
FPN-Fusion ICRA’22 (Tomy et al., 2022) $32.28\%$
EFNet ECCV’22 (Sun et al., 2022a) $35.33\%$
RENet ICRA’23 (Zhou et al., 2023) $38.38\%$
FlexEvent Ours - $\mathbf{48.64\%}$
Figure 5: Qualitative results of state-of-the-art event camera detectors. We compare FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors. See the Appendix for more examples.
Table 4: Ablation study of components in FlexEvent. FlexFuser denotes the adaptive event-frame fusion module (cf. Sec.3.2). FAL denotes the frequency-adaptive learning mechanism (cf. Sec. 3.3). The reported are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FAL.
Modality FAL FlexFuser Frequency (Hz) Average
$\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
E $53.2\%$ $54.0\%$ $53.5\%$ $52.0\%$ $49.4\%$ $45.9\%$ $38.8\%$ $22.9\%$ $46.2\%$
$54.6\%$ $54.9\%$ $54.9\%$ $54.3\%$ $53.3\%$ $50.7\%$ $44.6\%$ $30.4\%$ $49.7\%$
E + F $\blacklozenge$ $54.9\%$ $57.3\%$ $57.7\%$ $57.8\%$ $57.2\%$ $56.1\%$ $53.7\%$ $48.3\%$ $55.4\%$
$\mathbf{58.0\%}$ $59.6\%$ $\mathbf{60.0\%}$ $59.6\%$ $59.0\%$ $57.6\%$ $54.8\%$ $49.2\%$ $57.2\%$
$57.4\%$ $\mathbf{60.0\%}$ $\mathbf{60.0\%}$ $\mathbf{60.1\%}$ $\mathbf{59.5\%}$ $\mathbf{58.8\%}$ $\mathbf{56.5\%}$ $\mathbf{50.9\%}$ $\mathbf{57.9\%}$
Comparison to Multimodal Models. We compare FlexEvent with multimodal event-camera object detection methods such as DAGr (Gehrig & Scaramuzza, 2024), CAFR (Cao et al., 2024), and RENet (Zhou et al., 2023), which fuse event data with other sensor inputs, such as RGB frames or depth to improve detection accuracy. While these methods enhance performance over event-only approaches, they struggle to adapt to varying operational frequencies and often exhibit inadequate feature fusion in dynamic environments. Our approach addresses these limitations by dynamically balancing the contributions of event and frame data. As a result, we achieve superior performance, such as a $47.4\%$ mAP on DSEC-Detection as shown in Tab. 2, outperforming CAFR ($38.0\%$) and RENet ($29.4\%$). This flexible combination of event and frame data, along with its ability to generalize across different temporal resolutions, enables our method to excel in high-frequency detection scenarios, surpassing state-of-the-art methods.
Comparisons Across Different Categories. We evaluate the performance of FlexEvent across various object categories, including cars, pedestrians, and large vehicles. As shown in Tab. 2, we consistently outperform other methods, achieving $59.3\%$ mAP for cars, compared to $49.9\%$ for CAFR (Cao et al., 2024). This highlights its effectiveness in detecting larger, fast-moving objects, while also surpassing CAFR in pedestrian categories. Existing Event-based detectors struggle with detecting smaller and stationary objects due to a lack of pixel intensity changes, and fusion methods over-rely on frame data, which lacks temporal resolution. Our approach addresses these issues with adaptive fusion and temporal refinement, ensuring accurate detection across different object types and motion dynamics. This versatility reinforces its superiority over state-of-the-art methods.
Generalization on High-Frequency Data. A key contribution of FlexEvent is its ability to generalize across various operational frequencies, particularly in high-frequency scenarios. We evaluate this by testing detection performance at different temporal offsets, $\frac{i}{n}\Delta T$, where $n=10$, $i=0,...,10$, and $\Delta T=50$ ms. Ground truth labels are generated by linearly interpolating object positions between frames for consistent evaluation. In this setting, event-based methods are tested across multiple time durations, while event-frame fusion methods process one RGB frame followed by event data of varying time durations. The comparison results are shown in  Fig. 4. Most existing methods, such as RVT and SAST, struggle at higher frequencies due to their fixed temporal intervals and limited ability to capture fast scene changes. In contrast, our approach achieves $56.5\%$ mAP at $90$ Hz and $50.9\%$ at $180$ Hz. This improvement demonstrates that our method excels in dynamic, rapidly changing environments where accurate detection is critical for safety and reliability.
Qualitative Assessment. We provide qualitative comparisons between FlexEvent and other state-of-the-art methods under different event operation frequencies, as shown in Fig. 5 and Fig. 6. Unlike RVT and DAGr, which miss critical object details, our model consistently detects objects with high accuracy, even in challenging cases involving fast-moving vehicles and occluded pedestrians. For instance, in Scene 2 of Fig. 5, RVT fails to detect a pedestrian due to insufficient event data, while our approach successfully identifies the pedestrian by leveraging both frames and high-frequency event data. Similarly, in Scene 4, DAGr struggles with the rapid motion of a large vehicle, leading to inaccurate predictions, whereas our approach ensures precise object localization. In the upper instance of Fig. 6, RVT fails to detect a pedestrian at high frequency due to insufficient event data. In contrast, FlexEvent successfully identifies the pedestrian by leveraging the frame-based information. These qualitative findings confirm that our model excels not only in quantitative metrics but also in real-world performance.
Figure 6: Qualitative comparisons of FlexEvent and RVT (Gehrig & Scaramuzza, 2023) under different event operation frequencies. Our approach demonstrates a strong robustness under both low- and high-frequency scenarios. Best viewed in colors. See the Appendix for more examples.
Table 5: Ablation study of hyperparameter configurations in the FlexEvent frameworks. $\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively. $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length. The reported results are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
 $\mathbf{\tau}^{\text{car}}$  $\mathbf{\tau}^{\text{ped}}$ $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) Average
$\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
$0.6$ $0.3$ $10$ $0.8$ $56.5\%$ $55.9\%$ $56.7\%$ $57.2\%$ $57.1\%$ $56.7\%$ $54.5\%$ $49.2\%$ $55.5\%$
$0.6$ $0.3$ $10$ $0.6$ $56.7\%$ $57.2\%$ $57.7\%$ $57.9\%$ $57.7\%$ $57.0\%$ $54.3\%$ $47.0\%$ $55.7\%$
$0.6$ $0.3$ $8$ $0.6$ $56.3\%$ $58.5\%$ $58.8\%$ $59.1\%$ $58.8\%$ $58.4\%$ $56.2\%$ $\mathbf{51.2\%}$ $57.2\%$
$0.6$ $0.3$ $6$ $0.6$ $57.3\%$ $59.4\%$ $59.7\%$ $59.9\%$ $59.3\%$ $58.5\%$ $55.7\%$ $48.8\%$ $57.3\%$
$0.6$ $0.6$ $6$ $0.6$ $\mathbf{57.4\%}$ $\mathbf{60.0}\%$ $\mathbf{60.0\%}$ $\mathbf{60.1\%}$ $\mathbf{59.5\%}$ $\mathbf{58.8\%}$ $\mathbf{56.5\%}$ $50.9\%$ $\mathbf{57.9\%}$
$0.8$ $0.8$ $6$ $0.6$ $56.6\%$ $58.7\%$ $59.1\%$ $58.9\%$ $58.4\%$ $57.4\%$ $55.6\%$ $50.2\%$ $56.9\%$
4.3 Ablation studies
Component Analysis. We conduct ablation studies by selectively removing key modules: the FlexFuser and FAL mechanisms. As shown in Tab. 4, removing the FAL mechanism results in a significant performance drop, particularly in high-frequency scenarios, underscoring its role in adapting to varying frequencies and generating frequency-adjusted labels. Similarly, omitting the FlexFuser module leads to a marked decrease in mAP, highlighting the importance of adaptive event-frame fusion for accurate detection across different operational frequencies. We also test training with interpolated labels for high-frequency testing, but this approach reduces recall by missing objects that suddenly appear or disappear, making it less effective than FAL.
Hyperparameter Tuning. We tune the hyperparameters of the FAL mechanism, focusing on key settings like the confidence threshold ($\tau$), IoU threshold, and track length for temporal refinement. As shown in Tab. 5, lowering the confidence threshold improves recall but reduces precision, as the model becomes more lenient in detecting objects. Applying overly strict conditions, such as a higher confidence threshold or IoU threshold, lowers recall by filtering out valid detections. The optimal configuration is achieved with $\tau=0.6$ and a track length of 6, balancing precision and recall for both low- and high-frequency conditions. These moderate settings ensure that FlexEvent maintains robust performance and stable detection accuracy across diverse environments.
5 Conclusion
This paper introduces FlexEvent, an event-camera object detection framework designed to operate across arbitrary frequencies. By combining FlexFuser for adaptive event-frame fusion and FAL for frequency-adaptive learning, we leverage the rich temporal information of event data and the semantic detail of RGB frames to overcome the limitations of existing methods, providing a flexible solution for dynamic environments. Extensive experiments on large-scale datasets show that our approach significantly outperforms state-of-the-art methods, particularly in high-frequency scenarios, demonstrating its robustness and adaptability for real-world applications such as autonomous driving and robotics.
Appendix
In this appendix, we supplement the following materials to support the findings and conclusions drawn in the main body of this paper.
[…]
Appendix A A Additional Implementation Details
In this section, to facilitate future reproductions, we elaborate on the necessary details in terms of the datasets, evaluation metrics, and implementation details adopted in our experiments.
A.1 Datasets
In this work, we develop and validate our proposed approach on the large-scale DSEC dataset (Gehrig et al., 2021b). DSEC serves as a high-resolution, large-scale multimodal dataset designed to capture real-world driving scenarios under various conditions. It combines data from stereo Prophesee Gen3 event cameras with a resolution of $640\times 480$ pixels and FLIR Blackfly S RGB cameras operating at $20$ FPS, enabling high-fidelity capture of dynamic scenes. To align the RGB frames with the event camera data, an infinite-depth alignment process is employed, which involves undistorting, rotating, and re-distorting the RGB images. This alignment ensures that the event data and RGB frames are temporally and spatially synchronized.
Table 6: Summary of key statistics from the event camera object detection datasets used in this work.
Dataset Reference Classes Frames Sequences Class Names
DSEC-MOD (Zhou et al., 2023) $1$ $13,314$ $16$ Car
DSEC-Detection (Tomy et al., 2022) $3$ $52,727$ $41$ Car
Pedestrian
Large-Vehicle
DSEC-Det (Gehrig & Scaramuzza, 2024) $8$ $78,344$ $60$ Car
Pedestrian
Bus
Bicycle
Truck
Motorcycle
Rider
Train
In our experiments, we utilize three comprehensive versions of DSEC tailored for object detection: DSEC-Det (Gehrig & Scaramuzza, 2024), DSEC-Detection (Tomy et al., 2022), and DSEC-MOD (Zhou et al., 2023). A summary of the key statistics of these datasets is listed in Tab. 6.
  • DSEC-Det (Gehrig & Scaramuzza, 2024): This version is developed by the original DSEC team and includes annotations generated using the QDTrack multi-object tracker (Fischer et al., 2023; Pang et al., 2021). The annotation process combines automated multi-object tracking with manual refinement to ensure high-quality and accurate detection labels. Compared to the original DSEC dataset, DSEC-Det introduces additional sequences specifically designed to capture complex and dynamic urban environments, featuring crowded pedestrian areas, moving vehicles, and diverse lighting conditions. These challenging scenarios provide a rich testing ground for evaluating object detection algorithms in real-world driving settings. DSEC-Det comprises $60$ sequences and $78,344$ frames, making it the most extensive dataset used in this study. It captures diverse urban scenes with dynamic elements, such as crowded pedestrian areas and moving vehicles. Covering eight object categories relevant to autonomous driving – Car, Pedestrian, Bus, Bicycle, Truck, Motorcycle, Rider, and Train – the dataset provides a robust foundation for training and evaluating object detection models across various driving scenarios. In our experiments on DSEC-Det, to maintain consistency with the experimental setup of previous work DAGr (Gehrig & Scaramuzza, 2024), we report results on two categories: Car and Pedestrian.
  • DSEC-Detection (Tomy et al., 2022): This dataset comprises $41$ sequences with a total of $52,727$ frames. Focusing on three fundamental object categories – Car, Pedestrian, and Large Vehicle – this version emphasizes high-precision annotations for these critical classes in autonomous driving. The initial annotations are generated using the YOLOv5 model (Jocher, 2020) on RGB frames, leveraging its robust performance in real-time object detection. These annotations are then transferred to the corresponding event frames through homographic transformation, ensuring spatial alignment between the two modalities. A subsequent manual refinement process corrects any discrepancies and enhances annotation quality, resulting in a dataset that provides accurate and reliable labels for event-based object detection.
  • DSEC-MOD (Zhou et al., 2023): This dataset extends the object detection capabilities to moving-object detection across diverse urban environments. It includes $16$ sequences containing $13,314$ frames and is specifically focused on the Car category, making it highly suitable for complex detection tasks in varied urban settings, such as intersections, highways, and residential areas. Featuring high-frequency and dense annotations, the dataset provides a valuable resource for evaluating the performance of event-based object detectors under challenging real-world conditions.
These three versions of the DSEC dataset together provide a comprehensive platform for benchmarking and evaluating event-based object detection methods, capturing a wide spectrum of scenarios, object categories, and environmental conditions. Among them, DSEC-Det is the largest, most recent and most comprehensive, annotated and released by the original DSEC authors. Thus, we prioritize it as the primary benchmark for reporting results, ensuring relevance and reliability. DSEC-Detection and DSEC-MOD are datasets used by two recent event-frame fusion methods, CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023). To validate the effectiveness of our method, we also report results on these two datasets.
A.2 Baselines
To evaluate the effectiveness of our method, we compare it against both event-only and event-frame fusion state-of-the-art methods.
  • Event-Only Methods. We include state-of-the-art event-only object detectors, namely RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), LEOD (Wu et al., 2024), and SSM (Zubić et al., 2024), which are originally trained on event-only datasets like Gen1 (De Tournemire et al., 2020) and 1Mpx (Perot et al., 2020). To ensure a fair comparison, we retrain these methods on the DSEC-Det dataset following their respective training protocols.
  • Event-Frame Fusion Methods. For event-frame fusion methods on DSEC-Det, we include DAGr, as it has been evaluated on this dataset. We report the scores of DAGr (Gehrig & Scaramuzza, 2024) from the original paper to ensure consistency and fairness. For the DSEC-Detection and DSEC-MOD datasets, we train our model following the standard training and evaluation settings, and compare it against state-of-the-art methods CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023), as reported in their respective papers. For other methods evaluated on DSEC-Detection and DSEC-MOD, we reference the results reported in the CAFR and RENet papers, respectively. Since DAGr’s training code is not publicly available, we could not reproduce its results on DSEC-Detection and DSEC-MOD.
These comparisons ensure a fair and comprehensive evaluation while adhering to resource and code availability constraints.
A.3 Evaluation Metrics
In this work, we adopt the mean Average Precision (mAP) as the primary metric to evaluate the performance of our object detection models, consistent with standard practices in the field. The mAP metric provides a comprehensive measure of detection accuracy across multiple categories and intersection-over-union (IoU) thresholds.
Mathematically, the Average Precision (AP) for a single class is calculated as:
[…]
where $p(r)$ represents the precision at a given recall level $r$. The mean Average Precision (mAP) is then computed as the mean of the AP values across all object categories and a range of IoU thresholds (typically from $0.5$ to $0.95$ with a step size of $0.05$). This provides an overall measure of model performance across different levels of localization precision.
In addition to mAP, we also report the following metrics from the COCO evaluation protocol (Lin et al., 2014):
  • AP50: The average precision when evaluated at a fixed IoU threshold of 0.50, indicating how well the model performs with relatively lenient localization criteria.
  • AP75: The average precision at a fixed IoU threshold of 0.75, representing performance under stricter localization requirements.
  • APS, APM, and APL: These metrics represent the average precision for small ($S$), medium ($M$), and large ($L$) objects, respectively. Object sizes are defined based on their pixel area, with APS typically representing objects with areas less than $32\times 32$ pixels, APM representing areas between $32\times 32$ and $96\times 96$ pixels, and APL for objects larger than $96\times 96$ pixels.
By reporting these metrics, we obtain a more nuanced understanding of the model’s detection capabilities across varying object sizes and localization precision levels, ensuring a comprehensive evaluation of detection performance.
A.4 Training & Inference Details
We train our models using mixed precision to optimize both memory efficiency and training speed. The training process spans 100,000 iterations, utilizing the ADAM optimizer (Kingma, 2014) with a OneCycle learning rate schedule (Smith & Topin, 2019), which gradually decays from a peak learning rate to enhance convergence.
Consistent with RVT (Gehrig & Scaramuzza, 2023), we employ a mixed batching strategy to balance computational efficiency and memory usage. Specifically:
  • Standard Backpropagation Through Time (BPTT): Applied to half of the training samples, allowing for full sequence training.
  • Truncated BPTT (TBPTT): Used for the other half, reducing memory usage by splitting sequences into smaller segments.
For data augmentation, we apply random horizontal flipping and zoom transformations (both zoom-in and zoom-out) to enhance the diversity of training samples.
Our training process utilize the YOLOX framework (Zheng et al., 2021), a versatile object detection framework known for its efficient and high-performing architecture. We employ a multi-component loss function to optimize our model effectively:
  • Intersection over Union (IoU) Loss: This loss component measures the overlap between the predicted bounding boxes and the ground-truth boxes, ensuring that the predicted regions closely match the actual object locations.
  • Classification Loss: This component evaluates the accuracy of class predictions for each detected object, ensuring that the model correctly identifies the category of each detected instance.
  • Regression Loss: This loss assesses the precision of the predicted bounding box coordinates, helping the model refine the location and size of bounding boxes to align closely with the ground-truth annotations.
To ensure stable training, these loss components are averaged across both the batch and sequence length at each optimization step. This averaging process helps to reduce variance during training and facilitates smoother convergence of the model parameters.
Training Configuration. The training is conducted with a batch size of 8, which provides an optimal balance between efficient GPU utilization and memory requirements. Each training sample contains a sequence length of 11 frames, allowing the model to learn temporal dependencies effectively. The frame backbone’s weights are initialized using pre-trained ResNet and the event backbone’s weights are initialized using pre-trained RVT. The learning rate is set to $1\times 10^{-4}$, following a OneCycle learning rate schedule that allows efficient exploration of the learning space and helps in achieving faster convergence.
Hardware & Training Time. All training experiments are carried out on two NVIDIA RTX A5000 GPUs, each with 24GB of memory, providing the computational resources necessary for handling the high-resolution event data and RGB frames. The complete training process, including all iterations and model optimization, takes approximately one day, demonstrating the efficiency of our implementation in terms of both training speed and resource utilization.
Table 7: Comparative efficiency analysis of state-of-the-art event camera detectors on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024), comparing both event-only and event-frame fusion methods. This table reports inference times at various frequencies, measured in milliseconds (ms).
Modality Method Param Frequency (Hz)
(M) $\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30}$ $\mathbf{36}$ $\mathbf{45}$ $\mathbf{60}$ $\mathbf{90}$ $\mathbf{180}$ $\mathbf{200}$
E RVT $18.5$ $9.20$ $8.67$ $8.35$ $7.93$ $7.61$ $7.51$ $7.19$ $6.77$ $6.34$
SAST 18.9 $14.06$ $13.11$ $12.68$ $12.37$ $11.95$ $11.63$ $11.52$ $11.10$ $10.36$
SSM 18.2 $8.79$ $8.26$ $8.08$ $7.71$ $7.55$ $7.30$ $6.90$ $6.54$ $6.12$
E + F DAGr-50 $34.6$ $73.35$ $65.73$ $60.02$ $55.11$ $51.00$ $48.00$ $45.29$ $43.89$ $37.58$
FlexEvent $45.4$ $14.27$ $13.53$ $13.32$ $13.00$ $12.79$ $12.58$ $12.47$ $12.37$ $12.12$
Table 8: The complete results of the ablation study (cf. Tab. 4) of different components in the FlexEvent framework. FlexFuser denotes the adaptive event-frame fusion module (cf. Sec.3.2). FAL denotes the frequency-adaptive learning module (cf. Sec. 3.3). The reported results are the mAP, AP50, AP75, APS, APM, and APL scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FAL. The best and 2nd best scores of each metric are highlighted in bold and underline, respectively.
Modality FAL FlexFuser Frequency (Hz) mAP AP50 AP75 APS APM APL
Event $20.0$ $53.2\%$ $\mathbf{77.2\%}$ $58.1\%$ $\mathbf{46.4\%}$ $64.4\%$ $83.0\%$
$27.5$ $\mathbf{54.0\%}$ $76.8\%$ $\mathbf{59.3\%}$ $\mathbf{46.4\%}$ $66.6\%$ $\mathbf{85.2\%}$
$30.0$ $53.5\%$ $75.5\%$ $\mathbf{59.3\%}$ $45.6\%$ $\mathbf{66.8\%}$ $85.0\%$
$36.0$ $52.0\%$ $73.3\%$ $58.1\%$ $44.0\%$ $65.5\%$ $84.9\%$
$45.0$ $49.4\%$ $69.5\%$ $55.4\%$ $40.7\%$ $64.1\%$ $84.3\%$
$60.0$ $45.9\%$ $64.2\%$ $51.8\%$ $36.5\%$ $62.3\%$ $82.7\%$
$90.0$ $38.8\%$ $55.4\%$ $43.9\%$ $28.5\%$ $55.3\%$ $79.9\%$
$180.0$ $22.9\%$ $36.1\%$ $23.9\%$ $14.1\%$ $34.5\%$ $60.1\%$
Event $20.0$ $54.6\%$ $\mathbf{79.1\%}$ $\mathbf{61.8\%}$ $47.4\%$ $64.4\%$ $81.4\%$
$27.5$ $\mathbf{54.9\%}$ $78.8\%$ $61.4\%$ $\mathbf{47.6\%}$ $66.1\%$ $83.2\%$
$30.0$ $\mathbf{54.9\%}$ $78.5\%$ $61.3\%$ $47.4\%$ $\mathbf{66.9\%}$ $83.3\%$
$36.0$ $54.3\%$ $77.1\%$ $60.5\%$ $46.8\%$ $66.7\%$ $83.4\%$
$45.0$ $53.3\%$ $75.3\%$ $59.8\%$ $45.6\%$ $65.4\%$ $\mathbf{83.8\%}$
$60.0$ $50.7\%$ $72.4\%$ $57.3\%$ $42.3\%$ $63.5\%$ $83.5\%$
$90.0$ $44.6\%$ $65.1\%$ $49.9\%$ $35.3\%$ $58.9\%$ $81.9\%$
$180.0$ $30.4\%$ $48.1\%$ $32.2\%$ $20.7\%$ $44.0\%$ $72.9\%$
Event + Frame $\blacklozenge$ $20.0$ $54.9\%$ $74.0\%$ $63.2\%$ $50.7\%$ $61.3\%$ $85.5\%$
$\blacklozenge$ $27.5$ $57.3\%$ $75.7\%$ $66.3\%$ $\mathbf{52.8\%}$ $65.8\%$ $86.9\%$
$\blacklozenge$ $30.0$ $57.7\%$ $\mathbf{75.9\%}$ $\mathbf{66.8\%}$ $52.7\%$ $67.2\%$ $\mathbf{87.5\%}$
$\blacklozenge$ $36.0$ $\mathbf{57.8\%}$ $75.7\%$ $66.5\%$ $52.5\%$ $67.9\%$ $87.2\%$
$\blacklozenge$ $45.0$ $57.2\%$ $75.5\%$ $65.4\%$ $51.6\%$ $\mathbf{68.2\%}$ $\mathbf{87.5\%}$
$\blacklozenge$ $60.0$ $56.1\%$ $74.2\%$ $63.4\%$ $50.1\%$ $68.1\%$ $86.5\%$
$\blacklozenge$ $90.0$ $53.7\%$ $72.2\%$ $59.5\%$ $47.1\%$ $66.2\%$ $85.7\%$
$\blacklozenge$ $180.0$ $48.3\%$ $66.9\%$ $52.2\%$ $40.8\%$ $60.6\%$ $84.2\%$
Event + Frame $20.0$ $58.0\%$ $76.5\%$ $66.4\%$ $52.7\%$ $66.2\%$ $86.3\%$
$27.5$ $59.6\%$ $\mathbf{78.2\%}$ $\mathbf{69.6\%}$ $\mathbf{54.1\%}$ $69.9\%$ $\mathbf{88.0\%}$
$30.0$ $\mathbf{60.0\%}$ $78.1\%$ $69.5\%$ $53.7\%$ $\mathbf{71.3\%}$ $87.8\%$
$36.0$ $59.6\%$ $77.2\%$ $68.6\%$ $53.1\%$ $71.1\%$ $87.7\%$
$45.0$ $59.0\%$ $76.7\%$ $67.1\%$ $52.1\%$ $71.1\%$ $87.8\%$
$60.0$ $57.6\%$ $75.2\%$ $65.6\%$ $50.2\%$ $70.6\%$ $87.2\%$
$90.0$ $54.8\%$ $72.6\%$ $61.9\%$ $46.8\%$ $68.8\%$ $86.3\%$
$180.0$ $49.2\%$ $67.4\%$ $53.5\%$ $40.8\%$ $62.3\%$ $85.4\%$
Event + Frame $20.0$ $57.4\%$ $78.2\%$ $66.6\%$ $51.7\%$ $64.9\%$ $83.7\%$
$27.5$ $60.0\%$ $79.4\%$ $70.1\%$ $53.5\%$ $68.4\%$ $\mathbf{86.1\%}$
$30.0$ $60.0\%$ $\mathbf{79.7\%}$ $\mathbf{70.8\%}$ $\mathbf{53.6\%}$ $69.9\%$ $\mathbf{86.1\%}$
$36.0$ $\mathbf{60.1\%}$ $79.6\%$ $\mathbf{70.8\%}$ $53.2\%$ $70.3\%$ $85.7\%$
$45.0$ $59.5\%$ $79.0\%$ $69.5\%$ $52.5\%$ $70.8\%$ $85.3\%$
$60.0$ $58.8\%$ $78.5\%$ $69.0\%$ $51.1\%$ $\mathbf{71.1\%}$ $85.3\%$
$90.0$ $56.5\%$ $76.5\%$ $65.4\%$ $48.2\%$ $70.1\%$ $83.8\%$
$180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
Table 9: The complete results of the ablation study (cf. Tab. 5) of different hyperparameter configurations in the FlexEvent framework. $\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively. $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length. The reported results are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024). The best and 2nd best scores of each metric from each hyperparameter configuration are highlighted in bold and underline, respectively.
 $\mathbf{\tau}^{\text{car}}$  $\mathbf{\tau}^{\text{ped}}$  $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) mAP AP50 AP75 APS APM APL
$0.6$ $0.3$ $10$ $0.8$ $20.0$ $56.5\%$ $\mathbf{81.3\%}$ $66.4\%$ $51.8\%$ $62.2\%$ $82.8\%$
$0.6$ $0.3$ $10$ $0.8$ $27.5$ $55.9\%$ $74.7\%$ $65.1\%$ $51.9\%$ $61.4\%$ $87.0\%$
$0.6$ $0.3$ $10$ $0.8$ $30.0$ $56.7\%$ $75.3\%$ $66.0\%$ $52.3\%$ $63.5\%$ $87.0\%$
$0.6$ $0.3$ $10$ $0.8$ $36.0$ $\mathbf{57.2\%}$ $75.9\%$ $\mathbf{67.0\%}$ $\mathbf{52.5\%}$ $64.8\%$ $86.9\%$
$0.6$ $0.3$ $10$ $0.8$ $45.0$ $57.1\%$ $75.7\%$ $66.2\%$ $51.5\%$ $66.1\%$ $87.2\%$
$0.6$ $0.3$ $10$ $0.8$ $60.0$ $56.7\%$ $75.3\%$ $65.7\%$ $50.3\%$ $\mathbf{67.4\%}$ $\mathbf{87.3\%}$
$0.6$ $0.3$ $10$ $0.8$ $90.0$ $54.5\%$ $73.2\%$ $62.3\%$ $47.3\%$ $66.3\%$ $86.2\%$
$0.6$ $0.3$ $10$ $0.8$ $180.0$ $49.2\%$ $68.2\%$ $54.2\%$ $40.8\%$ $62.4\%$ $85.5\%$
$0.6$ $0.3$ $10$ $0.6$ $20.0$ $56.7\%$ $\mathbf{80.6\%}$ $65.5\%$ $51.2\%$ $63.0\%$ $81.7\%$
$0.6$ $0.3$ $10$ $0.6$ $27.5$ $57.2\%$ $79.3\%$ $65.0\%$ $51.9\%$ $65.2\%$ $84.5\%$
$0.6$ $0.3$ $10$ $0.6$ $30.0$ $57.7\%$ $79.4\%$ $66.0\%$ $\mathbf{52.3\%}$ $66.3\%$ $85.0\%$
$0.6$ $0.3$ $10$ $0.6$ $36.0$ $\mathbf{57.9\%}$ $79.5\%$ $\mathbf{66.4\%}$ $52.2\%$ $66.8\%$ $84.8\%$
$0.6$ $0.3$ $10$ $0.6$ $45.0$ $57.7\%$ $79.2\%$ $65.6\%$ $51.7\%$ $67.1\%$ $85.1\%$
$0.6$ $0.3$ $10$ $0.6$ $60.0$ $57.0\%$ $78.8\%$ $64.8\%$ $50.5\%$ $\mathbf{67.6\%}$ $\mathbf{85.3\%}$
$0.6$ $0.3$ $10$ $0.6$ $90.0$ $54.3\%$ $76.4\%$ $60.0\%$ $46.9\%$ $66.3\%$ $84.5\%$
$0.6$ $0.3$ $10$ $0.6$ $180.0$ $47.0\%$ $69.0\%$ $49.1\%$ $37.8\%$ $61.1\%$ $83.9\%$
$0.6$ $0.3$ $8$ $0.6$ $20.0$ $56.3\%$ $77.2\%$ $64.9\%$ $50.4\%$ $64.1\%$ $83.3\%$
$0.6$ $0.3$ $8$ $0.6$ $27.5$ $58.5\%$ $78.3\%$ $68.1\%$ $52.5\%$ $66.8\%$ $84.8\%$
$0.6$ $0.3$ $8$ $0.6$ $30.0$ $58.8\%$ $78.5\%$ $68.7\%$ $52.6\%$ $67.9\%$ $85.7\%$
$0.6$ $0.3$ $8$ $0.6$ $36.0$ $\mathbf{59.1\%}$ $\mathbf{78.8\%}$ $\mathbf{69.0\%}$ $\mathbf{52.7\%}$ $68.8\%$ $\mathbf{86.5\%}$
$0.6$ $0.3$ $8$ $0.6$ $45.0$ $58.8\%$ $78.2\%$ $68.6\%$ $52.3\%$ $68.8\%$ $86.1\%$
$0.6$ $0.3$ $8$ $0.6$ $60.0$ $58.4\%$ $77.9\%$ $67.5\%$ $51.3\%$ $\mathbf{69.6\%}$ $85.5\%$
$0.6$ $0.3$ $8$ $0.6$ $90.0$ $56.2\%$ $76.6\%$ $64.8\%$ $48.5\%$ $68.3\%$ $84.7\%$
$0.6$ $0.3$ $8$ $0.6$ $180.0$ $51.2\%$ $71.9\%$ $56.3\%$ $42.6\%$ $64.2\%$ $82.9\%$
$0.6$ $0.3$ $6$ $0.6$ $20.0$ $57.3\%$ $80.0\%$ $65.2\%$ $51.2\%$ $65.8\%$ $84.1\%$
$0.6$ $0.3$ $6$ $0.6$ $27.5$ $59.4\%$ $81.3\%$ $68.5\%$ $53.4\%$ $68.8\%$ $\mathbf{85.7\%}$
$0.6$ $0.3$ $6$ $0.6$ $30.0$ $59.7\%$ $\mathbf{81.7\%}$ $\mathbf{69.0\%}$ $\mathbf{53.7\%}$ $69.0\%$ $85.3\%$
$0.6$ $0.3$ $6$ $0.6$ $36.0$ $\mathbf{59.9\%}$ $81.4\%$ $\mathbf{69.0\%}$ $53.6\%$ $\mathbf{69.8\%}$ $\mathbf{85.7\%}$
$0.6$ $0.3$ $6$ $0.6$ $45.0$ $59.3\%$ $80.5\%$ $67.9\%$ $52.8\%$ $69.4\%$ $85.6\%$
$0.6$ $0.3$ $6$ $0.6$ $60.0$ $58.5\%$ $79.6\%$ $67.0\%$ $51.5\%$ $69.6\%$ $84.9\%$
$0.6$ $0.3$ $6$ $0.6$ $90.0$ $55.7\%$ $77.2\%$ $62.8\%$ $48.1\%$ $67.9\%$ $84.3\%$
$0.6$ $0.3$ $6$ $0.6$ $180.0$ $48.8\%$ $70.6\%$ $50.9\%$ $40.8\%$ $61.1\%$ $83.4\%$
$0.6$ $0.6$ $6$ $0.6$ $20.0$ $57.4\%$ $78.2\%$ $66.6\%$ $51.7\%$ $64.9\%$ $83.7\%$
$0.6$ $0.6$ $6$ $0.6$ $27.5$ $60.0\%$ $79.4\%$ $70.1\%$ $53.5\%$ $68.4\%$ $\mathbf{86.1\%}$
$0.6$ $0.6$ $6$ $0.6$ $30.0$ $60.0\%$ $\mathbf{79.7\%}$ $\mathbf{70.8\%}$ $\mathbf{53.6\%}$ $69.9\%$ $\mathbf{86.1\%}$
$0.6$ $0.6$ $6$ $0.6$ $36.0$ $\mathbf{60.1\%}$ $79.6\%$ $\mathbf{70.8\%}$ $53.2\%$ $70.3\%$ $85.7\%$
$0.6$ $0.6$ $6$ $0.6$ $45.0$ $59.5\%$ $79.0\%$ $69.5\%$ $52.5\%$ $70.8\%$ $85.3\%$
$0.6$ $0.6$ $6$ $0.6$ $60.0$ $58.8\%$ $78.5\%$ $69.0\%$ $51.1\%$ $\mathbf{71.1\%}$ $85.3\%$
$0.6$ $0.6$ $6$ $0.6$ $90.0$ $56.5\%$ $76.5\%$ $65.4\%$ $48.2\%$ $70.1\%$ $83.8\%$
$0.6$ $0.6$ $6$ $0.6$ $180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
$0.8$ $0.8$ $6$ $0.6$ $20.0$ $56.6\%$ $80.7\%$ $65.5\%$ $50.8\%$ $65.4\%$ $82.6\%$
$0.8$ $0.8$ $6$ $0.6$ $27.5$ $58.7\%$ $81.9\%$ $68.9\%$ $\mathbf{52.8\%}$ $68.6\%$ $84.9\%$
$0.8$ $0.8$ $6$ $0.6$ $30.0$ $\mathbf{59.1\%}$ $\mathbf{82.0\%}$ $\mathbf{69.2\%}$ $52.7\%$ $69.5\%$ $84.8\%$
$0.8$ $0.8$ $6$ $0.6$ $36.0$ $58.9\%$ $81.7\%$ $68.8\%$ $52.6\%$ $\mathbf{69.9\%}$ $85.0\%$
$0.8$ $0.8$ $6$ $0.6$ $45.0$ $58.4\%$ $81.4\%$ $67.7\%$ $51.6\%$ $69.8\%$ $\mathbf{85.1\%}$
$0.8$ $0.8$ $6$ $0.6$ $60.0$ $57.4\%$ $80.1\%$ $66.6\%$ $50.2\%$ $\mathbf{69.9\%}$ $84.3\%$
$0.8$ $0.8$ $6$ $0.6$ $90.0$ $55.7\%$ $78.6\%$ $63.7\%$ $47.8\%$ $68.9\%$ $84.2\%$
$0.8$ $0.8$ $6$ $0.6$ $180.0$ $50.2\%$ $74.0\%$ $55.2\%$ $41.5\%$ $63.7\%$ $83.1\%$
Appendix B B Additional Quantitative Results
B.1 Comparison of Efficiency
We present a comparative analysis of inference times and parameter counts for the evaluated methods in Tab. 7. All experiments are conducted on an NVIDIA RTX A5000 24GB GPU paired with an AMD EPYC 9354P 32-Core Processor operating at 3.8 GHz. The results demonstrate that, despite having a higher parameter count, FlexEvent achieves inference times comparable to the event-only method SAST and significantly outperforms the event-frame fusion method DAGr in terms of speed. Additionally, FlexEvent consistently surpasses DAGr in inference time across all tested frequencies. These results underscore the efficiency and rapid performance of FlexEvent, highlighting its suitability for real-time applications.
B.2 Complete Results of Ablation Study
We include the complete results of the ablation study in Tab. 8.
B.3 Complete Results of Hyperparameter Searching
We include the complete results of the hyperparameter searching in Tab. 9.
Appendix C C Additional Qualitative Results
C.1 Visual Comparisons of Event Camera Detectors
We include additional qualitative assessments in Fig. 7, Fig. 8, and Fig. 9.
C.2 Visual Comparisons under Different Frequencies
We include additional qualitative assessments in Fig. 10, Fig. 11, and Fig. 12.
Appendix D D Potential Societal Impact & Limitations
In this section, we discuss the potential societal impact of FlexEvent, including its positive contributions, broader implications, and known limitations. While our method offers significant advancements in event camera object detection, it is important to consider its broader consequences and areas for future improvement.
D.1 Societal Impact
The development of FlexEvent introduces several positive societal benefits, particularly in safety-critical applications such as autonomous driving, robotics, and surveillance. By enhancing the ability to detect fast-moving objects in real time, our framework can improve the responsiveness and safety of autonomous systems operating in dynamic environments. This is especially important for avoiding collisions or responding to hazards in high-speed scenarios. For example, autonomous vehicles equipped with our approach can better detect pedestrians, cyclists, and other vehicles in real time, potentially reducing accidents and saving lives.
Additionally, the computational efficiency provided by the adaptive event-frame fusion (FlexFuser) and frequency-adaptive learning (FAL) mechanisms reduces the need for resource-intensive training processes. This contributes to the broader societal goal of making advanced AI technologies more accessible and less energy-intensive, thereby minimizing the environmental impact of large-scale AI models. Our approach could also benefit industries beyond transportation, such as robotics for healthcare, industrial automation, and public safety.
D.2 Broader Impact
The broader implications of FlexEvent include its potential to advance the field of event-based vision and enable new applications where high temporal resolution is crucial. By overcoming the limitations of conventional fixed-frequency object detection methods, our approach paves the way for more flexible, adaptable AI systems. This could lead to improvements in areas such as drone navigation, real-time video analysis for security purposes, and human-robot collaboration, where detecting fast-moving objects and adapting to changing environments are critical.
Moreover, the development of efficient and scalable detection systems like our approach can drive further innovation in resource-constrained environments, such as low-power edge devices. These advancements could make high-performance detection systems more widely available, particularly in developing regions or areas with limited access to computational resources.
However, as with any powerful technology, there is a risk of misuse. Enhanced object detection capabilities could potentially be exploited for surveillance purposes, raising privacy concerns. As event camera technology becomes more widespread, it is important to establish ethical guidelines and regulatory frameworks to ensure that these systems are used responsibly, particularly when monitoring public spaces or collecting sensitive data.
D.3 Known Limitations
While FlexEvent demonstrates significant performance improvements, there are several known limitations to our approach.
Dependence on Event Camera Quality. The effectiveness of our approach relies on the quality of the event camera sensor. Inconsistent or noisy event data, especially under poor lighting or extreme weather conditions, could affect detection performance. Future work could explore robustness to sensor noise and adaptation to diverse environmental conditions.
Limited Generalization to Unseen Scenarios. Although our approach shows strong performance across varying frequencies, it may still face challenges in completely unseen environments, where the motion dynamics and scene conditions differ significantly from the training data. Investigating methods for domain adaptation or online learning could help improve generalization to new contexts.
Resource Requirements for High-Frequency Data. While FlexFuser mitigates the computational cost of training on high-frequency event data, processing extremely high-frequency event streams still requires substantial computational resources during inference. This could limit the scalability on resource-constrained devices or in real-time applications with stringent latency requirements.
Figure 7: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors.
Figure 8: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors.
Figure 9: Additional qualitative results of state-of-the-art event camera detectors. We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det. Best viewed in colors.
Figure 10: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Figure 11: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Figure 12: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario. The experiments are conducted on the test set of DSEC-Det. Best viewed in colors.
Appendix E E Public Resources Used
In this section, we acknowledge the public resources used, during the course of this work.
E.1 Public Datasets Used
E.2 Public Implementations Used

### v3

@@
-Dongyue Lu1, Lingdong Kong1, Gim Hee Lee1, Camille Simon Chane2, Wei Tsang Ooi1
+FlexEvent: Towards Flexible Event-Frame Object Detection at Varying Operational Frequencies
 Abstract
-Event cameras offer unparalleled advantages for real-time perception in dynamic environments, thanks to their microsecond-level temporal resolution and asynchronous operation.
-Existing event-based object detection methods, however, are limited by fixed-frequency paradigms and fail to fully exploit the high-temporal resolution and adaptability of event cameras.
-To address these limitations, we propose FlexEvent, a novel event camera object detection framework that enables detection at arbitrary frequencies.
-Our approach consists of two key components: FlexFuser, an adaptive event-frame fusion module that integrates high-frequency event data with rich semantic information from RGB frames, and FAL, a frequency-adaptive learning mechanism that generates frequency-adjusted labels to enhance model generalization across varying operational frequencies.
+Event cameras offer unparalleled advantages for real-time perception in dynamic environments, thanks to the microsecond-level temporal resolution and asynchronous operation.
+Existing event detectors, however, are limited by fixed-frequency paradigms and fail to fully exploit the high-temporal resolution and adaptability of event data.
+To address these limitations, we propose FlexEvent, a novel framework that enables detection at varying frequencies.
+Our approach consists of two key components: FlexFuse, an adaptive event-frame fusion module that integrates high-frequency event data with rich semantic information from RGB frames, and FlexTune, a frequency-adaptive fine-tuning mechanism that generates frequency-adjusted labels to enhance model generalization across varying operational frequencies.
 This combination allows our method to detect objects with high accuracy in both fast-moving and static scenarios, while adapting to dynamic environments.
@@
 Our framework sets a new benchmark for event-based object detection and paves the way for more adaptable, real-time vision systems.
-Figure 1: Event camera detection at varying frequencies.
-The performance of the classic RVT detector (Gehrig & Scaramuzza, 2023) drops significantly at higher event operational frequencies.
-Motivated by this, we propose FlexEvent, a robust and flexible detector that maintain high accuracy across a wide range of frequencies, ensuring strong adaptability in dynamic sensing environments.
+Figure 1: Illustrative examples of event camera object detection at varying frequencies.
+The detection performance of the classic RVT detector [18] tends to drop significantly at higher event operational frequencies.
+Motivated by this observation, we propose FlexEvent, a robust and flexible event-frame detector that maintains high detection accuracy across a wide range of frequencies (low, middle, high), ensuring strong adaptability in real-world, dynamic sensing environments.
 1 Introduction
-Event cameras have garnered significant attention for their ability to capture dynamic scenes with microsecond-level temporal resolution (Gallego et al., 2022).
-Unlike conventional RGB cameras, which capture entire frames at fixed intervals, event cameras operate asynchronously, responding to changes in pixel intensity at each location (Zou et al., 2022).
-This low-latency operation reduces motion blur and enables highly energy-efficient sensing, making event cameras ideal for real-time applications such as autonomous driving, robotics, and surveillance (Steffen et al., 2019; Kong et al., 2024b; Chen et al., 2023).
-Despite their potential, existing event camera object detection methods often fail to fully leverage the high-frequency temporal information captured by event cameras (Cordone et al., 2022; Gehrig & Scaramuzza, 2022; Jeziorek et al., 2023).
-Most approaches align event data with the lower frequency of RGB cameras by adopting a fixed time interval between event streams and frame-based annotations (Perot et al., 2020; Gehrig et al., 2021b).
-While this strategy simplifies data processing, it inevitably overlooks the rich temporal details embedded in high-frequency event streams.
-Given that human annotations are often synchronized with slower frame rates, current detection models miss valuable information from high-frequency event data, resulting in suboptimal performance when rapid object detection is required in dynamic environments (Messikommer et al., 2020; Schaefer et al., 2022).
-To address these limitations, we introduce FlexEvent, a novel event camera object detection framework designed to tackle the challenging problem of object detection at varying operational frequencies.
+Event cameras have garnered significant attention for their ability to capture dynamic scenes with microsecond-level temporal resolution [12].
+Unlike RGB cameras, which capture entire frames at fixed intervals, event cameras operate asynchronously, responding to changes in pixel intensity at each location [68].
+This low-latency operation reduces motion blur and enables highly energy-efficient sensing, making event cameras ideal for real-time applications [45, 30, 6].
+Despite their potential, existing event detectors often fail to fully leverage the high-frequency temporal information [8, 15, 23].
+Most approaches align event data with the lower frequency of frames by adopting fixed time intervals between event streams and frame-based annotations [40, 17].
+While this simplifies data processing, it inevitably overlooks the rich temporal details embedded in high-frequency event streams.
+Given that human annotations are often synchronized with slower frame rates, current detectors miss valuable information from higher frequencies, resulting in suboptimal performance when rapid object detection is required in dynamic environments [36, 43].
+To address these limitations, we introduce FlexEvent, a novel framework designed to tackle the challenging problem of object detection at varying operational frequencies.
 Our approach addresses the need for high-frequency detection in fast-changing environments, while adapting to different operational frequencies.
-We propose two key innovations: (1) FlexFuser, an adaptive event-frame fusion module, and (2) FAL, a frequency-adaptive learning mechanism.
-Flexible Event-Frame Fusion.
-The first component, FlexFuser, addresses the limitations of event data, which often lacks semantic and texture-rich information, especially at higher frequencies (Zhou et al., 2023), by integrating the rich spatial and semantic information from RGB frames with the high-temporal resolution of event streams.
+We propose two key innovations: 1) an adaptive event-frame fusion module, and 2) a frequency-adaptive fine-tuning mechanism.
+The first component, FlexFuse, addresses the limitations of event data, which often lacks semantic and texture-rich information, especially at higher frequencies [66], by integrating the rich spatial and semantic information from RGB frames with the high-temporal resolution of event streams.
 It enables high detection accuracy even in fast-moving environments.
 Furthermore, training on high-frequency event data is computationally expensive and impractical due to the significant human effort required to label such data.
-FlexFuser mitigates this by sampling event data at varying frequencies, aligning them with the normal frame rate during training, thus maintaining efficiency while preserving the high-frequency benefits at inference time.
-Frequency-Adaptive Learning.
-The second component, FAL, enhances the generalization capability of event camera detectors across varying operational frequencies, by generating frequency-adjusted labels for the unlabeled high-frequency data.
+FlexFuse mitigates this by sampling event data at varying frequencies, aligning them with the normal frame rate during training, thus maintaining efficiency while preserving the high-frequency benefits at inference time.
+The second component, FlexTune, enhances the generalization capability of event camera detectors across varying operational frequencies, by generating frequency-adjusted labels for the unlabeled high-frequency data.
 These labels allow the model to learn from high-frequency event streams without manual annotations, and iterative refinement through self-training ensures that the model remains robust across different motion dynamics and frequency settings.
@@
 Our approach consistently outperforms recent detectors across both standard and high-frequency settings.
-In particular, we achieve mAP gains of $\mathbf{15.5\%}$, $\mathbf{9.4\%}$, and $\mathbf{10.3\%}$ over the previous best-performing detectors on the DSEC-Det (Gehrig & Scaramuzza, 2024), DSEC-Detection (Tomy et al., 2022), and DSEC-MOD (Zhou et al., 2023) datasets, respectively.
+In particular, we achieve mAP gains of $\mathbf{15.5\%}$, $\mathbf{9.4\%}$, and $\mathbf{10.3\%}$ over the previous best-performing detectors on DSEC-Det [16], DSEC-Detection [50], and DSEC-MOD [66], respectively.
 Our model also maintains $\mathbf{96.2\%}$ of its performance when the operational frequency shifts from $\mathbf{20}$ Hz to $\mathbf{90}$ Hz, and delivers accurate detection at frequencies as high as $\mathbf{180}$ Hz, proving its robustness under extreme conditions.
-[…]
-[…]
-[…]
-[…]
+In summary, our contributions are listed as follows:
+$\bullet$ The proposed FlexEvent framework is aimed at tackling the challenging problem of event camera object detection at varying frequencies, being one of the first works to address this challenging problem explicitly in real-world conditions.
+$\bullet$ We propose FlexFuse, an adaptive event-frame fusion module that leverages the strengths of both event and frame modalities, enabling more efficient and accurate event-based object detection in dynamic environments.
+$\bullet$ We introduce FlexTune, a frequency-adaptive fine-tuning mechanism that generates frequency-adjusted labels and improves generalization across various motion frequencies.
+$\bullet$ We demonstrate that our approach achieves promising results across large-scale datasets, particularly in high-frequency scenarios, validating its effectiveness to handle real-world, safety-critical problems.
 2 Related Work
 Event Camera Object Detection.
-Event-based detection methods can be broadly split into two approaches: GNNs/SNNs and dense feed-forward models.
-GNNs build dynamic spatio-temporal graphs by subsampling events (Gehrig & Scaramuzza, 2022; Sun & Ji, 2023; Messikommer et al., 2020; Schaefer et al., 2022), but they face challenges in propagating information over large spatio-temporal regions, especially for slow-moving objects.
-SNNs offer efficient sparse information transmission but are often hindered by their non-differentiable nature, complicating optimization processes (Cuadrado et al., 2023; Cordone et al., 2022; Zhang et al., 2022).
+Event-based detectors can be broadly split into two approaches: GNNs/SNNs and dense feed-forward models.
+GNNs build dynamic spatio-temporal graphs by sub-sampling events [15, 47, 36, 43], but they face challenges in propagating information over large spatio-temporal regions, especially for slow-moving objects.
+SNNs offer efficient sparse information transmission but are often hindered by their non-differentiable nature, complicating optimization processes [9, 8, 59].
 Dense, feed-forward models represent the second approach.
-Initial methods using fixed temporal windows (Chen, 2018; Iacono et al., 2018; Jiang et al., 2019) struggle with slow-moving or stationary objects due to their limited capability to capture long-term temporal data.
-Subsequent advancements incorporate RNNs and transformers to enhance temporal modeling capabilities (Perot et al., 2020; Zubić et al., 2023; Li et al., 2022; Gehrig & Scaramuzza, 2023; Peng et al., 2024), but these models often still lack semantic richness and face difficulties in adapting to variable frequencies and highly dynamic scenarios.
+Initial methods using fixed temporal windows [5, 22, 25, 19] struggle with slow-moving or stationary objects due to their limited capability to capture long-term temporal data.
+Subsequent work incorporates RNNs and transformers to enhance temporal modeling [40, 69, 34, 18, 39], but these models still lack semantic richness and face difficulties in adapting to variable frequencies and highly dynamic scenarios.
 Event-Frame Multimodal Learning.
-To overcome the limited texture in event streams, multimodal fusion techniques combining event-based and frame-based data have gained traction across tasks, such as deblurring (Sun et al., 2022a; Zhang et al., 2020), depth estimation (Gehrig et al., 2021a; Uddin et al., 2022), and tracking (Zhao et al., 2022; Gehrig et al., 2020).
-Earlier object detection approaches fuse event and image data during post-processing (Li et al., 2019; Chen et al., 2019), but they lack meaningful feature-level interaction.
-Recent works focus on deeper feature fusion (Tomy et al., 2022; Cao et al., 2022; 2021), with advanced methods introducing pixel-level spatial attention or temporal transformers for asynchronous processing (Zhou et al., 2023; Li et al., 2023; Gehrig & Scaramuzza, 2024; Cao et al., 2024).
-Some approaches (Li et al., 2023; Gehrig et al., 2021a) explore combining events and frames through asynchronous multi-modal fusion, enabling inference at varying frequencies.
-However, they do not focus on high-frequency event data and fully leverage the temporally rich nature of event streams.
-Furthermore, these methods still face challenges in fully exploiting complementary strengths and addressing feature imbalance in event-frame detection.
-Unlike previous methods, FlexEvent employs a more comprehensive fusion framework that effectively combines high-temporal resolution event data with rich semantics from RGB frames, enabling robust object detection across varying frequencies while addressing feature imbalance.
-Label-Efficient Learning in Event Data.
+Combining events with frame-based data has proven effective in tasks such as deblurring [48, 61], depth estimation [14, 51], and tracking [62, 13].
+Early fusion methods perform post-processing combinations [33, 4], but lack feature-level interaction.
+More recent works propose deeper feature fusion techniques [50, 1, 2], introducing pixel-level attention or temporal transformers for asynchronous processing [66, 32, 16, 3].
+Some explore asynchronous multi-modal fusion for flexible inference at different frequencies [32, 14, 60], yet they do not explicitly address high-frequency event data or fully exploit temporal richness.
+Additionally, balancing contributions from event and frame modalities remains challenging.
+Unlike these works, FlexEvent introduces a unified fusion framework that adaptively integrates high-frequency event data with semantic-rich frames, achieving robust detection across diverse frequencies while addressing feature imbalance.
+Label-Efficient Learning from Event Data.
 Due to limited annotated datasets, label-efficient learning has become an important area for event-based vision.
-Several studies attempt to reconstruct images from event data (Rebecq et al., 2019; 2021; Stoffregen et al., 2020) or leverage knowledge distillation from pre-trained frame-based models (Wang et al., 2021; Sun et al., 2022b; Yang et al., 2023; Kong et al., 2024a).
-Other approaches utilize pre-trained models or self-supervised losses (Klenk et al., 2022; Wu et al., 2023; Zhu et al., 2019).
-LEOD (Wu et al., 2024) pioneers object detection with limited labels but does not address high-frequency generalization.
-A recent state-space model (Zubić et al., 2024) adapts to varying frequencies without retraining but struggles to detect static objects at high frequencies due to its reliance solely on event data.
+Several studies attempt to reconstruct images from event data [41, 42, 46] or distill knowledge from pre-trained frame-based models [52, 49, 57, 29].
+Other approaches use pre-trained models or self-supervised losses [28, 56, 67].
+LEOD [55] pioneers object detection with limited labels but does not address high-frequency generalization.
+A recent state-space model [70] adapts to varying frequencies without retraining but struggles to detect static objects at high frequencies due to its reliance solely on event data.
 In contrast, FlexEvent is specifically designed to adapt to varying event frequencies, ensuring consistent performance even in scenarios with limited labels, and effectively detecting both stationary and fast-moving objects.
-3 FlexEvent: A Flexible Event Object Detector
+3 FlexEvent: A Flexible Event-Frame Object Detector
 In this section, we elaborate on the technical details of our FlexEvent framework.
@@
 3.1.
-We then introduce the FlexFuser module in Sec.
+We then introduce the FlexFuse module in Sec.
 3.2, which adaptively fuses event and frame data to enhance detection across varying frequencies.
-Finally, we detail the frequency-adaptive learning (FAL) mechanism in Sec.
-3.3, which enables our model to generalize effectively across diverse temporal conditions using self-training and adaptive label generation.
-The overall framework is illustrated in Fig.2.
+Finally, we detail FlexTune, our frequency-adaptive fine-tuning mechanism, in Sec.
+3.3, which enables our model to generalize effectively across diverse temporal conditions via adaptive label generation.
+The overall framework is illustrated in Fig.
+2.
 Figure 2: Framework Overview.
@@
 3.1).
-These branches are fused dynamically through FlexFuser, allowing adaptive integration of event and frame data (cf.
+These branches are fused dynamically through FlexFuse, allowing adaptive integration of event and frame data (cf.
 Sec.
 3.2).
-Additionally, the frequency-adaptive learning (FAL) mechanism ensures robust detection performance across varying operational frequencies (cf.
+Additionally, the FlexTune learning mechanism ensures robust detection performance across varying operational frequencies (cf.
 Sec.
@@
 Event cameras are bio-inspired vision sensors that capture changes in log intensity per pixel asynchronously, rather than capturing entire frames at fixed intervals.
-Formally, let $I(x,y,t)$ denotes the log intensity at pixel coordinates $(x,y)$ and time $t$.
-An event $e$ is generated at $(x,y,t)$ whenever the change in log intensity $\Delta I$ exceeds a certain threshold $C$.
-Such a process can be modeled as:
+Formally, let $I(x,y,t)$ denote the log intensity at pixel coordinates $(x,y)$ and time $t$.
+An event $e$ is generated at $(x,y,t)$ whenever the change in log intensity $\Delta I$ exceeds a certain threshold $C$, which is modeled as:
 $\Delta I(x,y,t)=I(x,y,t)-I(x,y,t-\Delta t)\geq C~.$ (1)
 Each event $e$ is a tuple $(x,y,t,p)$, where $(x,y)$ are the pixel coordinates, $t$ is the timestamp, and $p\in\{-1,1\}$ denotes the polarity of the event which indicates the direction of the intensity change.
-To leverage event data with convolutional neural network layers, we preprocess events into a 4D tensor $E$ with dimensions representing the polarity, temporal discretization $T$, and spatial dimensions $(H,W)$.
-This representation involves mapping a set of events $\mathcal{E}$ within time interval $[t_{a},t_{b})$ into:
-$E(p,\tau,x,y)=\sum_{e_{k}\in\mathcal{E}}\delta(p-p_{k})\delta(x-x_{k},y-y_{k})\delta(\tau-\tau_{k}),$ (2)
-where $\tau_{k}=\left\lfloor\frac{t_{k}-t_{a}}{t_{b}-t_{a}}\cdot T\right\rfloor$.
-This tensor captures event activity in $T$ discrete time slices, yielding a compact representation suitable for 2D convolutions by flattening the polarity and temporal dimensions.
+To leverage event data with convolutions, we pre-process events into a 4D tensor $E$ with dimensions [18] representing the polarity, temporal discretization $T$, and spatial dimensions $(H,W)$, respectively.
+This representation involves mapping a set of events $\mathcal{E}$ within time interval $t_{1},t_{2})$ into the following:
+$E(p,\tau,x,y)=\sum\nolimits_{e_{k}\in\mathcal{E}}\delta(p-p_{k})\delta(x-x_{k},y-y_{k})\delta(\tau-\tau_{k}),$ (2)
+where $\tau_{k}=\left\lfloor\frac{t_{k}-t_{1}}{t_{2}-t_{1}}\cdot T\right\rfloor$.
+Such a 4D tensor captures event activity in $T$ discrete time slices, yielding a compact representation suitable for 2D operators by flattening the polarity and temporal dimensions.
 Problem Formulation.
-Given two consecutive frames $F_{a}$ and $F_{b}$ captured at timestamps $T_{a}$ and $T_{b}$, our objective is to leverage the event stream over the interval $[T_{a},T_{b}]$ to detect objects at the end timestamp $T_{b}$.
-Existing event-based object detection methods often use fixed time intervals $\Delta T$, limiting adaptability to dynamic environments (Perot et al., 2020).
-Additionally, integrating semantic information from RGB frames remains challenging, affecting performance in complex scenarios (De Tournemire et al., 2020).
-To address this, we synchronize the event data with frames and explore varying training frequencies, leveraging the temporal richness of event cameras to improve detection accuracy.
-3.2 FlexFuser: An Adaptive Event-Frame Fusion Module
-In dynamic environments, object detection systems must adapt to varying motion frequencies (Sun et al., 2022a).
-While event cameras excel at capturing rapid changes in pixel intensity, they often lack the rich spatial and semantic information provided by frames.
-To address this limitation and fully leverage the complementary strengths of both modalities, we introduce FlexFuser, an adaptive fusion module designed to dynamically combine event data at different frequencies with frame data.
+Given two consecutive frames $F_{1}$ and $F_{2}$ captured at timestamps $t_{1}$ and $t_{2}$, our objective is to leverage the event stream over the interval $t_{1},t_{2}]$ to detect objects at the end timestamp $t_{2}$.
+Existing event detectors use fixed time intervals $\Delta t$, limiting adaptability to dynamic environments [40].
+Additionally, integrating semantic information from RGB frames remains challenging, affecting performance in complex scenarios [10].
+To address this, we propose to synchronize the event data with frames and explore varying training frequencies, leveraging the temporal richness of event cameras to improve detection accuracy.
+3.2 FlexFuse: Adaptive Event-Frame Fusion
+In dynamic environments, object detection systems must adapt to varying motion frequencies [48].
+While event cameras excel at capturing rapid changes in pixel intensity, they often lack rich spatial and semantic information provided by frames.
+To address this limitation and fully leverage the complementary strengths of both modalities, we introduce FlexFuse, an adaptive fusion module designed to dynamically combine event data at different frequencies with frames.
 Dynamic Event Aggregation.
-Given a dataset $\mathcal{D}$, consisting of sequences of calibrated event camera data and frame data with a resolution of $H\times W$, along with corresponding bounding box annotations $y$ collected at frequency $a$, we begin by selecting a batch of frame data $\mathbf{F}$ paired with event data $\mathbf{E}^{a}$, both captured at frequency $a$.
+Given a dataset $\mathcal{D}$, consisting of sequences of calibrated event data and frame data with a resolution of $H\times W$, along with corresponding bounding box annotations $y$ collected at frequency $a$, we begin by selecting a batch of frame data $\mathbf{F}$ paired with event data $\mathbf{E}^{a}$, both captured at frequency $a$.
 To aggregate event data from a higher frequency $b$ (where $b>a$), we divide the time interval $\Delta t^{a}$ corresponding to $\mathbf{E}^{a}$ into $b/a$ smaller sub-intervals.
@@
 2.
-From this set, we randomly sample one event data point [^1] For simplicity, we use $\mathbf{E}^{b}$ to represent a sample from the set of high-frequency event data $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, rather than explicitly referencing each individual sample from the event set.
+From this set, we randomly sample one data point [^1] For simplicity, we use $\mathbf{E}^{b}$ to represent a sample from the set of high-frequency event data $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, rather than explicitly referencing each individual sample from the event set.
 The same applies to other frequencies.
 $\mathbf{E}^{b}$.
-By doing so, the sampled high-frequency event data $\mathbf{E}^{b}$ is temporally aligned with the frame data $\mathbf{F}$ and the base frequency event data $\mathbf{E}^{a}$.
-This synchronization of event streams at different frequencies ensures consistent and reliable processing for subsequent stages.
-Figure 3: Illustration of the FlexFuser module. We show a general example of event and frame under frequency $a$ at stage $i$.
+As detections are predicted for the end of $\Delta t^{a}$, this strategy pairs each frame with the preceding high‐frequency events while introducing only millisecond‐scale jitter that acts as implicit temporal augmentation, strengthening robustness to real‐world synchronization noise.
+Consequently, $(\mathbf{F},\mathbf{E}^{a},\mathbf{E}^{b})$ are effectively paired across modalities and frequencies, and this synchronization of image streams with event streams at different frequencies ensures consistent and reliable processing for all subsequent stages.
 Feature Extraction.
-Let $\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}(\cdot)$ and $\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}(\cdot)$ represent the event- and frame-based networks, respectively, where the former employs the RVT (Gehrig & Scaramuzza, 2023) for extracting features from event data, and the latter uses ResNet-50 (He et al., 2016) for feature extraction from frames.
+Let $\phi_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}(\cdot)$ and $\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}(\cdot)$ represent the event- and frame-based networks, respectively, where the former employs the RVT [18] for extracting features from event data, and the latter uses ResNet-50 [20] for feature extraction from frames.
 Both networks are structured into four stages, as shown in Fig.
 2.
-At each scale $i$, we extract features ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a},^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}$ from the event data and ${}^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}$ from the frame data:
-${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}=\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{(i)}(\mathbf{E}^{a}),\quad^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}=\phi_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{(i)}(\mathbf{E}^{b}),\quad^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}=\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}^{(i)}(\mathbf{F}),$ (3)
-where ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}$ and ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{b}\in\mathbb{R}^{B\times C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}\times H_{i}\times W_{i}}$, ${}^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\in\mathbb{R}^{B\times C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\times H_{i}\times W_{i}}$.
-Here, $i$ denotes the scale, $B$ is the batch size, and $C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}$ and $C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}$ are the dimensions of the feature maps extracted from the event and frame data, respectively.
-Event-Frame Adaptive Fuser.
-To effectively fuse the event and frame data, we employ an adaptive fuser that is consistent across different event data frequencies.
-At each scale $i$, taking the low-frequency event features ${}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}$ as an example, we concatenate the feature maps from both the event and frame branches as follows:
-${}^{(i)}\mathbf{h}_{\text{shared}}^{a}=\left[{}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a},\ ^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right]\in\mathbb{R}^{B\times\left(C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right)\times H_{i}\times W_{i}}~{}.$ (4)
-Inspired by previous works (Zhou et al., 2023) and (Zhong et al., 2024), our goal is to dynamically fuse these two modalities in a flexible manner.
-The proposed FlexFuser module computes adaptive soft weights that regulate the contribution of each branch (event and frame) based on the current input conditions.
+At each scale $i$, we extract features $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a},^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{b}$ from the event data and $^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}$ from the frame data.
+This process is formulated as:
+$^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}=^{(i)}\phi_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}(\mathbf{E}^{a}),^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{b}=^{(i)}\phi_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}(\mathbf{E}^{b}),^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}=^{(i)}\phi_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}(\mathbf{F}),$ (3)
+where $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}$ and $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{b}\in\mathbb{R}^{B\times C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}\times H_{i}\times W_{i}}$, $^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\in\mathbb{R}^{B\times C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\times H_{i}\times W_{i}}$.
+Here, $i$ denotes the scale, $B$ is the batch size, and $C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}$ and $C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}$ are the dimensions of the event and frame feature maps, respectively.
+Figure 3: Illustration of the FlexFuse module. We show a general example of event and frame under frequency $a$ at the $i$-stage.
+Event-Frame Adaptive Fusion.
+To effectively fuse the event and frame data, we employ an adaptive fusion that is consistent across different event data frequencies.
+At each scale $i$, taking the low-frequency event features $^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}$ as an example, we concatenate the feature maps from both the event and frame branches as follows:
+$^{(i)}\mathbf{h}_{\mathrm{shared}}^{a}=\left[{}^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a},\ ^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\right]\in\mathbb{R}^{B\times\left(C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\right)\times H_{i}\times W_{i}}~.$ (4)
+Inspired by previous work [66, 64], our goal is to dynamically fuse these two modalities in a flexible manner.
+The proposed FlexFuse module computes adaptive soft weights that regulate the contribution of each branch (event and frame) based on the current input conditions.
 As shown in Fig.
-3, these adaptive soft weights are computed using a gating function, which incorporates learned noise to introduce perturbations for improved adaptability.
-The process is:
-$\left[{}^{(i)}\bm{\alpha},\ ^{(i)}\bm{\beta}\right]=\text{Softmax}\left((^{(i)}\mathbf{h}_{\text{shared}}^{a}\cdot^{(i)}\mathbf{W})+^{(i)}\sigma\cdot\epsilon\right),$ (5)
-where ${}^{(i)}\mathbf{W}\in\mathbb{R}^{\left(C_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}}\right)\times 2}$ is a trainable weight matrix, ${}^{(i)}\bm{\alpha}$ and ${}^{(i)}\bm{\beta}$ are the adaptive soft weights for the event and frame branches, respectively.
+3, these adaptive soft weights are computed using a gating function, which incorporates learned noise to introduce perturbations for improved adaptability:
+$^{(i)}\boldsymbol{\alpha},\ ^{(i)}\boldsymbol{\beta}]=\mathrm{Softmax}((^{(i)}\mathbf{h}_{\mathrm{shared}}^{a}\cdot^{(i)}\mathbf{W})+^{(i)}\sigma\cdot\epsilon),$ (5)
+where $^{(i)}\mathbf{W}\in\mathbb{R}^{\left(C_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}+C_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}}\right)\times 2}$ is a trainable weight matrix, $^{(i)}\boldsymbol{\alpha}$ and $^{(i)}\boldsymbol{\beta}$ are the adaptive soft weights for the event and frame branches, respectively.
 Here, $^{(i)}\sigma$ is a learned standard deviation that controls the magnitude of the noise perturbation, and $\epsilon\sim\mathcal{N}(0,1)$ represents a Gaussian noise term.
-The fused feature map at each scale $i$ is then obtained by applying the adaptive soft weights to the event and frame features:
-${}^{(i)}\mathbf{h}_{\text{fuse}}^{a}=^{(i)}\bm{\alpha}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.70703125,0.546875}\text{E}}}^{a}+^{(i)}\bm{\beta}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\definecolor[named]{pgfstrokecolor}{rgb}{0.94140625,0.38671875,0.29296875}\text{F}}},$ (6)
+The fused feature map at each scale $i$ is then obtained by applying adaptive soft weights:
+$^{(i)}\mathbf{h}_{\mathrm{fuse}}^{a}=^{(i)}\boldsymbol{\alpha}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0,0.70703125,0.546875}\mathrm{E}}}^{a}+^{(i)}\boldsymbol{\beta}\odot^{(i)}\mathbf{h}_{{\color[rgb]{0.94140625,0.38671875,0.29296875}\mathrm{F}}},$ (6)
 where $\odot$ denotes element-wise multiplication.
-This fusion process dynamically balances the contribution of each modality based on the input data, allowing for more robust and adaptive feature representation across varying conditions.
+This fusion process dynamically balances the contribution of each modality based on the input data, allowing for more adaptive feature representation across varying conditions.
 Then, at each scale $i$, the final feature map combining event data at different frequencies and the frame data is obtained by adding the fused features from different frequencies.
-Specifically, we combine the fused feature maps as ${}^{(i)}\mathbf{h}_{\text{fuse}}=^{(i)}\mathbf{h}_{\text{fuse}}^{a}+^{(i)}\mathbf{h}_{\text{fuse}}^{b}$.
+Specifically, we combine the fused feature maps as $^{(i)}\mathbf{h}_{\mathrm{fuse}}=^{(i)}\mathbf{h}_{\mathrm{fuse}}^{a}+^{(i)}\mathbf{h}_{\mathrm{fuse}}^{b}$.
 After obtaining the fused feature maps across all scales, the multi-scale features are concatenated and fed into the detection head to produce the predicted bounding box $\mathbf{\hat{y}}$.
-Optimization & Regularization.
-In addition to the standard detection loss $\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}})$, we introduce a regularization term to ensure balanced utilization of both the event and frame branches.
-This term penalizes large variations in the soft weights, encouraging a more uniform contribution from both modalities and preventing overfitting to a single branch:
-$\mathcal{L}_{\text{fuser}}=\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}})+\lambda\left(\frac{\text{Var}\left(\bm{\alpha}\right)}{\left(\mathbb{E}\left[\bm{\alpha}\right]\right)^{2}}+\frac{\text{Var}\left(\bm{\beta}\right)}{\left(\mathbb{E}\left[\bm{\beta}\right]\right)^{2}}\right),$ (7)
+[…]
+$\mathcal{L}_{\mathrm{fuse}}=\mathcal{L}_{\mathrm{det}}(\mathbf{y},\mathbf{\hat{y}})+\lambda\left(\frac{\mathrm{Var}\left(\boldsymbol{\alpha}\right)}{\left(\mathbb{E}\left[\boldsymbol{\alpha}\right]\right)^{2}}+\frac{\mathrm{Var}\left(\boldsymbol{\beta}\right)}{\left(\mathbb{E}\left[\boldsymbol{\beta}\right]\right)^{2}}\right),$ (7)
 where $\lambda$ is a weighting factor.
-3.3 FAL: Frequency-Adaptive Learning Mechanism
-FlexFuser aggregates information from different frequencies using labeled low-frequency data.
-To tune the model adaptively to handle diverse frequencies by leveraging both labeled low-frequency data and unlabeled high-frequency data, we design a flexible frequency-adaptive learning (FAL) mechanism.
-FAL incorporates multi-frequency information into the training process through iterative self-training.
-This approach enhances the model’s ability to generalize across varying frequencies, making it more robust in different scenarios.
-The key steps of the FAL mechanism are as follows:
-Pre-Training with Low-Frequency Labels.
-Rather than training solely at the data collection frequency $a$, we enhance the model’s capability by training it at a higher frequency $b$.
-To efficiently leverage the available labels, we select only the final event from the high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, which corresponds to the labeled timestamp.
-This approach allows the model to capture valuable high-frequency temporal information while still utilizing low-frequency labels, improving its temporal understanding and robustness.
-The training objective is to minimize the detection loss over the labeled data as follows:
-$\mathcal{L}_{\text{GT}}=\sum_{(\mathbf{F},\mathbf{E}^{b}_{b/a},\mathbf{y})\in\mathcal{D}}\mathcal{L}_{\text{det}}(\mathbf{y},\mathbf{\hat{y}}).$ (8)
-Labels at Higher Frequencies.
-For the unlabeled data in $\mathcal{D}$ captured at frequency $b$, the pre-trained model generates high-frequency labels $\mathbf{\hat{y}}$ by performing inference on the entire high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$.
-These generated labels $\mathbf{\hat{y}}$ serve as pseudo-labels for guiding further training at higher frequencies, improving the model’s ability to generalize across different temporal conditions.
-Enhanced Temporal Refinement.
-To refine the high-frequency labels, we introduce a multi-step temporal refinement approach.
-First, we adopt bidirectional event augmentation by processing both forward and reversed event streams to detect objects with varying movements and orientations, thereby boosting recall.
-After generating the bidirectional labels, we apply Non-Maximum Suppression (NMS) to remove overlapping bounding boxes, followed by a low confidence threshold $\mathbf{\tau}$ to retain potential objects and further improve recall.
-Next, leveraging a tracking-by-detection framework, we link detection boxes across frames using pairwise IoU matching with a threshold $\mathbf{\tau}^{\text{IoU}}$.
-Short-lived tracks, with lengths below $\mathbf{L}^{\text{track}}$, are pruned to ensure temporal consistency.
-This approach ensures that the refined high-frequency labels $\tilde{\mathbf{y}}$ are accurate, temporally consistent, and reliable, ultimately improving detection quality in high-frequency data even in the absence of ground truth labels.
-Self-Training Iteration.
-The model is iteratively trained using the refined high-frequency labels $\tilde{\mathbf{y}}$ on high-frequency data where no ground truth labels are available.
+3.3 FlexTune: Frequency-Adaptive Fine-Tuning
+FlexFuse aggregates information from different frequencies using labeled low-frequency data.
+To tune the model adaptively to handle diverse frequencies by leveraging both labeled low-frequency data and unlabeled high-frequency data, we design a FlexTune learning mechanism, incorporating multi-frequency information into the training process through iterative self-training.
+This enhances the ability to generalize across varying frequencies, making it more robust in different scenarios.
+As shown in Fig.
+4, FlexTune consists of two main stages: Low-Frequency Sparse Training to learn foundational knowledge from labeled data, and Cross-Frequency Propagation to transfer and refine the learned knowledge with high-frequency unlabeled data.
+This iterative mechanism bridges sparse supervision with dense temporal patterns, enabling robust detection under varying frequencies.
+Figure 4: Illustration of the FlexTune learning mechanism.
+We first train on high-frequency events with sparse low-frequency labels, then we generate and refine high-frequency labels for cyclic self-training across frequencies.
+Low-Frequency Sparse Training.
+Rather than training solely at frequency $a$, we enhance the model’s capability by training it at a higher frequency $b$.
+To efficiently leverage the available sparse labels, we select only the last event from the high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$, which corresponds to the labeled timestamp.
+This allows the model to capture valuable high-frequency temporal information while still utilizing low-frequency labels, improving its temporal understanding and robustness.
+The training objective is to minimize the detection loss over the sparse labeled data as:
+[…]
+where the summation is taken over samples $(\mathbf{F},\mathbf{E}^{b}_{b/a},\mathbf{y})\in\mathcal{D}$.
+Cross-Frequency Propagation.
+Building on the low-frequency training, we transfer and refine model knowledge to unlabeled high-frequency data through three sequential steps: High-Frequency Bootstrapping, Temporal Consistency Calibration, and Cyclic Self-Training.
+Together, these steps generate accurate high-frequency pseudo-labels, enabling robust detection in diverse temporal settings.
+High-Frequency Bootstrapping.
+For the unlabeled data in $\mathcal{D}$ captured at frequency $b$, the pre-trained model generates high-frequency labels $\mathbf{\tilde{y}}$ by performing inference on the entire high-frequency event set $\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a}$.
+These generated labels $\mathbf{\tilde{y}}$ serve as pseudo-labels for bootstrapping further training at higher frequencies, improving the model’s ability to generalize across different temporal conditions.
+Temporal Consistency Calibration.
+To refine high-frequency labels, we enforce temporal coherence through three steps.
+(1) Bidirectional Event Augmentation.
+Process event streams forward and backward to capture diverse object motions, enhancing recall.
+(2) Confidence-Aware Filtering.
+Apply Non-Maximum Suppression (NMS) and a low confidence threshold $\mathbf{\tau}$ to eliminate duplicates and retain high-potential detections.
+(3) Tracklet Pruning.
+Link detections across frames using IoU-based tracking $(\mathbf{\tau}^{\mathrm{IoU}})$ and prune short tracks $(<\mathbf{L}^{\mathrm{track}})$ to suppress transient noise.
+This approach ensures that the refined high-frequency labels $\tilde{\mathbf{y}}$ are accurate, temporally consistent, and reliable, improving detection quality in high-frequency data even in the absence of ground truth labels.
+Cyclic Self-Training.
+The model is iteratively trained using the refined high-frequency labels $\tilde{\mathbf{y}}$ and the ground truth low-frequency label.
 The total loss function combines the base training loss and the pseudo-label loss as:
-$\mathcal{L}_{\text{FAL}}=\mathcal{L}_{\text{GT}}+\beta\sum_{(\mathbf{F},\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a-1},\mathbf{\tilde{y}})\in\mathcal{D}}\mathcal{L}_{\text{det}}\left(\mathbf{\tilde{y}},\mathbf{\hat{y}}\right)~{},$ (9)
-where $\beta$ balances the contribution of the high-frequency label loss.
-The complete FlexEvent framework combines FlexFuser and FAL, allowing the model to dynamically fuse event and frame data while adapting to varying frequencies.
-As demonstrated in the following sections, this combination provides a robust detection framework capable of maintaining high accuracy in dynamic environments.
+[…]
+where the summation here is taken over high-frequency samples $(\mathbf{F},\{\mathbf{E}_{i}^{b}\}_{i=0}^{b/a-1},\mathbf{\tilde{y}})\in\mathcal{D}$, and the coefficient $\beta$ balances the contribution of the high-frequency label loss.
+The complete FlexEvent framework combines FlexFuse and FlexTune, allowing the model to dynamically fuse event and frame data while adapting to varying frequencies.
 4 Experiments
-Table 1: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
+4.1 Experimental Settings
+Datasets.
+We conduct experiments based on three large-scale datasets: 1DSEC-Det [16], 2DSEC-Detection [50], and 3DSEC-MOD [66].
+These datasets comprise 78,34478{,}344 frames across $60$ sequences, 52,72752{,}727 frames over $41$ sequences, and 13,31413{,}314 frames within $16$ sequences, respectively.
+making them suitable for evaluating event-based object detection methods.
+We prioritize DSEC-Det [16] as the primary benchmark for comparisons, as it is the largest, most recent, and most comprehensive event-frame perception dataset.
+We report results on Car and Pedestrian classes to ensure fair comparison with the state-of-the-art method DAGr [16].
+For more details, please refer to Appendix Sec.
+A.1.
+Table 1: Comparative study of state-of-the-art event camera detectors on the validation set of the DSEC-Det [16] dataset.
 Both event-only and event-frame fusion methods are compared.
 The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
-Modality Method Venue Reference mAP AP50 AP75 APS APM APL
-E RVT CVPR’23 (Gehrig & Scaramuzza, 2023) $38.4\%$ $58.7\%$ $41.3\%$ $29.5\%$ $50.3\%$ $81.7\%$
-SAST CVPR’24 (Peng et al., 2024) $38.1\%$ $60.1\%$ $40.0\%$ $29.8\%$ $48.9\%$ $79.7\%$
-SSM CVPR’24 (Zubić et al., 2024) $38.0\%$ $55.2\%$ $40.6\%$ $28.8\%$ $52.2\%$ $77.8\%$
-LEOD CVPR’24 (Wu et al., 2024) $41.1\%$ $65.2\%$ $43.6\%$ $35.1\%$ $47.3\%$ $73.3\%$
-E + F DAGr-18 Nature’24 (Gehrig & Scaramuzza, 2024) $37.6\%$
-DAGr-34 Nature’24 (Gehrig & Scaramuzza, 2024) $39.0\%$
-DAGr-50 Nature’24 (Gehrig & Scaramuzza, 2024) $41.9\%$ $66.0\%$ $44.3\%$ $36.3\%$ $56.2\%$ $77.8\%$
-FlexEvent Ours - $\mathbf{57.4\%}$ $\mathbf{78.2\%}$ $\mathbf{66.6\%}$ $\mathbf{51.7\%}$ $\mathbf{64.9\%}$ $\mathbf{83.7\%}$
-4.1 Experimental Settings
-Datasets.
-We conduct experiments based on three large-scale event camera datasets: 1DSEC-Det (Gehrig & Scaramuzza, 2024), 2DSEC-Detection (Tomy et al., 2022), and 3DSEC-MOD (Zhou et al., 2023).
-These datasets comprise $78,344$ frames across $60$ sequences, $52,727$ frames over $41$ sequences, and $13,314$ frames within $16$ sequences, respectively, making them suitable for evaluating event-based object detection methods.
-We prioritize DSEC-Det as the primary benchmark for reporting results as it is the largest, most recent, and most comprehensive event-frame perception dataset.
-For more details, please refer to the Appendix.
+   Modality Method Ref Venue mAP AP50 AP75 APS APM APL
+E RVT [18] CVPR’23 $38.4$ $58.7$ $41.3$ $29.5$ $50.3$ $81.7$
+SAST [39] CVPR’24 $38.1$ $60.1$ $40.0$ $29.8$ $48.9$ $79.7$
+SSM [70] CVPR’24 $38.0$ $55.2$ $40.6$ $28.8$ $52.2$ $77.8$
+LEOD [55] CVPR’24 $41.1$ $65.2$ $43.6$ $35.1$ $47.3$ $73.3$
+E + F HDI-Former [31] arXiv’24 $46.7$ $69.1$ - - - -
+DAGr-18 [16] Nature’24 $37.6$ - - - - -
+DAGr-34 [16] Nature’24 $39.0$ - - - - -
+     DAGr-50 [16]    Nature’24 $41.9$ $66.0$ $44.3$ $36.3$ $56.2$ $77.8$
+FlexEvent - Ours $\mathbf{57.4}$ $\mathbf{78.2}$ $\mathbf{66.6}$ $\mathbf{51.7}$ $\mathbf{64.9}$ $\mathbf{83.7}$
+Table 2: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Detection [50].
+Both event-only and event-frame fusion methods are compared.
+The reported results are the mAP scores of the 1Car, 2Pedestrian, and 3Large-Vehicle (L-Veh.) classes.
+The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
+   Modality Method Venue Ref Type   Car Pedestrian  L-Veh. Average
+E CAFR    ECCV’24 [3] Event - - - $12.0$
+SENet CVPR’18 [21] Attention  $38.4$ $14.9$ $26.0$ $26.2$
+CBAM ECCV’18 [54]  $37.7$ $13.5$ $27.0$ $26.1$
+ECA-Net CVPR’20 [53]  $36.7$ $12.8$ $27.5$ $25.7$
+SAGate ECCV’20 [7] RGB + Depth  $32.5$ $10.4$ $16.0$ $19.6$
+DCF CVPR’21 [24]  $36.3$ $12.7$ $28.0$ $25.7$
+SPNet ICCV’21 [65]  $39.2$ $17.8$ $26.2$ $27.7$
+RAMNet RA-L’21 [14] $24.4$ $10.8$ $17.6$ $17.6$
+FAGC Sensors’21 [2] $39.8$ $14.4$ $33.6$ $29.3$
+E + F    FPN-Fusion ICRA’22 [50]    RGB + Event  $37.5$ $10.9$ $24.9$ $24.4$
+EFNet ECCV’22 [48] $41.1$ $15.8$ $32.6$ $30.0$
+DRFuser EAAI’23 [37] $38.6$ $15.1$ $30.6$ $28.1$
+CMX TITS’23 [58]  $41.6$ $16.4$ $29.4$ $29.1$
+RENet ICRA’23 [66]  $40.5$ $17.2$ $30.6$ $29.4$
+CAFR ECCV’24 [3]  $49.9$ $25.8$ $38.2$ $38.0$
+FlexEvent Ours -  $\mathbf{59.3}$ $\mathbf{37.4}$ $\mathbf{45.5}$ $\mathbf{47.4}$
+Figure 5: Qualitative comparisons of state-of-the-art event camera detectors. We compare FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det. Best viewed in colors.
+Figure 6: Results of detectors under varied frequencies on DSEC-Det.
+Table 3: Comparisons of state-of-the-art event-frame fusion detectors on the test set of the DSEC-MOD [66] dataset.
+Modal Method Venue Type F.mAP V.mAP Avg
+E+F SENet [21] CVPR’18 Attention $29.3$ $17.5$ $23.4$
+CBAM [54] ECCV’18 $36.2$ $16.4$ $26.3$
+ECA-Net [53] CVPR’20 $34.5$ $18.8$ $26.7$
+SAGate [7] ECCV’20 RGB + Depth $33.6$ $17.8$ $25.7$
+DCF [24] CVPR’21 $32.2$ $18.8$ $25.5$
+SPNet [65] ICCV’21 $32.7$ $14.7$ $23.7$
+FPN-Fusion [50] ICRA’22 RGB + Event $32.3$ $15.0$ $23.7$
+EFNet [48] ECCV’22 $35.3$ $18.1$ $26.7$
+RENet [66] ICRA’23 $38.4$ $19.6$ $29.0$
+FlexEvent Ours $\mathbf{48.6}$ $\mathbf{25.2}$ $\mathbf{36.9}$
+Table 4: Comparative efficiency analysis of event detectors on DSEC-Det [16]. We report the inference times of event detectors at various frequencies, measured in milliseconds (ms).
+Method Size Operational Frequency
+(M) $\mathbf{20.0}$ Hz $\mathbf{36.0}$ Hz $\mathbf{90.0}$ Hz $\mathbf{180}$ Hz
+RVT [18] $18.5$ $9.20$ ms $7.93$ ms $7.19$ ms $6.77$ ms
+SAST [39] $18.9$ $14.06$ ms $12.37$ ms $11.52$ ms $11.10$ ms
+SSM [70] $18.2$ $8.79$ ms $7.71$ ms $6.90$ ms $6.54$ ms
+DAGr-50 [16] $34.6$ $73.35$ ms $55.11$ ms $45.29$ ms $43.89$ ms
+FlexEvent $45.4$ $14.27$ ms $13.00$ ms $12.47$ ms $12.37$ ms
+Table 5: Ablation on FlexFuse and FlexTune learning mechanism on DSEC-Det [16]. Symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FlexTune.
+Tune Fuse Frequency (Hz) Avg
+$\mathbf{20.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
+      $53.2$ $52.0$ $49.4$ $45.9$ $38.8$ $22.9$ $43.7$
+$54.6$ $54.3$ $53.3$ $50.7$ $44.6$ $30.4$ $48.0$
+$\blacklozenge$ $54.9$ $57.8$ $57.2$ $56.1$ $53.7$ $48.3$ $54.7$
+$\mathbf{58.0}$ $59.6$ $59.0$ $57.6$ $54.8$ $49.2$ $56.4$
+$57.4$ $\mathbf{60.1}$ $\mathbf{59.5}$ $\mathbf{58.8}$ $\mathbf{56.5}$ $\mathbf{50.9}$ $\mathbf{57.2}$
 Baselines.
-To evaluate our method, we compare it against both event-only and event-frame fusion state-of-the-art methods.
-For event-only detectors, we include RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), LEOD (Wu et al., 2024), and SSM (Zubić et al., 2024), retraining them on the DSEC-Det dataset following their respective protocols.
-For event-frame fusion methods, we compare with DAGr (Gehrig & Scaramuzza, 2024) on DSEC-Det using results from the original paper and retrain our method on DSEC-Detection and DSEC-MOD using standard settings to compare with CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023).
+To evaluate our method, we compare it against both state-of-the-art event-only and event-frame fusion methods.
+For event-only detectors, we compare RVT [18], SAST [39], LEOD [55], and SSM [70], retraining them on DSEC-Det [16] following their respective protocols .
+For event-frame fusion methods, we compare DAGr [16] and HDI-Former [31] on DSEC-Det [16] using results from the original paper and retrain our method on DSEC-Detection [50] and DSEC-MOD [66] using standard settings to compare with CAFR [3] and RENet [66].
 For other methods on DSEC-Detection and DSEC-MOD, we reference results reported in the CAFR and RENet papers.
-For more details, please refer to the Appendix.
+For more details, please refer to Appendix Sec.
+A.2.
+Evaluation Metrics.
+We evaluate object detectors using the mean Average Precision (mAP) as the primary metric, along with AP50, AP75, APS, APM, and APL from the COCO evaluation protocol [35].
+These metrics provide a comprehensive assessment of detection performance across different IoU thresholds and object sizes.
+Please refer to the Appendix Sec.
+A.3 for more details.
 Implementation Details.
-We train our model using the YOLOX framework (Zheng et al., 2021), optimizing with the standard detection loss, which consists of IoU loss, classification loss, and regression loss, averaged across both the batch and sequence length for stable training.
-Additionally, we introduce an extra regularization term to balance the utilization among the different modalities.
-The model is trained for $100,000$ iterations with a batch size of $8$ and a sequence length of $11$, using a learning rate of $1\times 10^{-4}$.
-All experiments are conducted on two NVIDIA RTX A5000 GPUs with 24GB memory, with the entire training process completed in approximately one day.
-Due to space limits, more details are included in the Appendix.
-Evaluation Metrics.
-We evaluate object detectors using the mean Average Precision (mAP) as the primary metric, along with AP50, AP75, APS, APM, and APL from the COCO evaluation protocol (Lin et al., 2014).
-These metrics provide a comprehensive assessment of detection performance across different IoU thresholds and object sizes.
-Please refer to the Appendix for more details.
-4.2 Comparisons to State-of-the-Art Detectors
-Comparison to Event-Only Models.
-We compare FlexEvent with state-of-the-art event-only detectors, including RVT (Gehrig & Scaramuzza, 2023), SSM (Zubić et al., 2024), SAST (Peng et al., 2024), and LEOD (Wu et al., 2024), as shown in Tab.
+Our training follows YOLO-X [63], using the standard detection loss consisting of IoU loss, classification loss, and regression loss, plus a lightweight regularizer that balances the contributions of each modality.
+The model is trained for 100,000100{,}000 iterations with a batch size of $8$ and a sequence length of $11$, using a learning rate of 11e-44.
+Experiments are conducted on two NVIDIA RTX A50005000 GPUs with 2424GB memory, with the entire training process completed in approximately one day.
+Due to space limits, more details are in Appendix Sec.
+A.4.
+4.2 Comparisons to State of the Arts
+Comparisons with Event-Only Detectors.
+We compare FlexEvent with state-of-the-art event-only detectors, including RVT [18], SSM [70], SAST [39], and LEOD [55], as shown in Tab.
 1.
-We significantly outperform these methods across all metrics, with the performance gap becoming even more pronounced at higher frequencies.
-Event-only methods struggle to maintain detection accuracy at higher frequencies due to their inability to fully capture object semantics.
-In contrast, we overcome these limitations through the FlexFuser module, which integrates RGB data to compensate for the lack of semantic richness in the event stream.
+Our model substantially outperforms all these methods, with the gap widening at higher event frequencies.
+Event-only methods struggle to maintain detection accuracy at higher frequencies because they lack rich semantic cues.
+In contrast, we overcome these limitations through the FlexFuse module, which integrates RGB data to compensate for the lack of semantic richness in the event stream.
 By fusing both event and frame data, we excel in complex and dynamic environments, achieving superior detection accuracy where event-only methods fall short.
-Table 2: Comparative study of state-of-the-art event camera detectors on the test set of DSEC-Detection (Tomy et al., 2022).
-Both event-only and event-frame fusion methods are compared.
-The reported results are the mAP scores of 1Car, 2Pedestrian (Ped), and 3Large-Vehicle (L-Veh) classes.
-The best and 2nd best scores from each metric are highlighted in bold and underlined, respectively.
-Modality Method Venue Reference Type Car Ped L-Veh Average
-E CAFR ECCV’24 (Cao et al., 2024) Event - - - $12.0\%$
-E + F SENet CVPR’18 (Hu et al., 2018) Attention $38.4\%$ $14.9\%$ $26.0\%$ $26.2\%$
-CBAM ECCV’18 (Woo et al., 2018) $37.7\%$ $13.5\%$ $27.0\%$ $26.1\%$
-ECA-Net CVPR’20 (Wang et al., 2020) $36.7\%$ $12.8\%$ $27.5\%$ $25.7\%$
-SAGate ECCV’20 (Chen et al., 2020) RGB + Depth $32.5\%$ $10.4\%$ $16.0\%$ $19.6\%$
-DCF CVPR’21 (Ji et al., 2021) $36.3\%$ $12.7\%$ $28.0\%$ $25.7\%$
-SPNet ICCV’21 (Zhou et al., 2021) $39.2\%$ $17.8\%$ $26.2\%$ $27.7\%$
-RAMNet RA-L’21 (Gehrig et al., 2021a) $24.4\%$ $10.8\%$ $17.6\%$ $17.6\%$
-FAGC Sensors’21 (Cao et al., 2021) $39.8\%$ $14.4\%$ $33.6\%$ $29.3\%$
-FPN-Fusion ICRA’22 (Tomy et al., 2022) RGB + Event $37.5\%$ $10.9\%$ $24.9\%$ $24.4\%$
-EFNet ECCV’22 (Sun et al., 2022a) $41.1\%$ $15.8\%$ $32.6\%$ $30.0\%$
-DRFuser EAAI’23 (Munir et al., 2023) $38.6\%$ $15.1\%$ $30.6\%$ $28.1\%$
-CMX TITS’23 (Zhang et al., 2023) $41.6\%$ $16.4\%$ $29.4\%$ $29.1\%$
-RENet ICRA’23 (Zhou et al., 2023) $40.5\%$ $17.2\%$ $30.6\%$ $29.4\%$
-CAFR ECCV’24 (Cao et al., 2024) $49.9\%$ $25.8\%$ $38.2\%$ $38.0\%$
-FlexEvent Ours - $\mathbf{59.3\%}$ $\mathbf{37.4\%}$ $\mathbf{45.5\%}$ $\mathbf{47.4\%}$
-Figure 4: Comparisons of event-based detectors under different event frequencies on DSEC-Det.
-Table 3: Comparisons of fusion-based event detectors on the test set of DSEC-MOD (Zhou et al., 2023). The best and 2nd best scores from each metric are highlighted in bold and underlined.
-Method Venue Reference mAP
-SENet CVPR’18 (Hu et al., 2018) $29.28\%$
-CBAM ECCV’18 (Woo et al., 2018) $36.22\%$
-ECA-Net CVPR’20 (Wang et al., 2020) $34.49\%$
-SAGate ECCV’20 (Chen et al., 2020) $33.62\%$
-DCF CVPR’21 (Ji et al., 2021) $32.20\%$
-SPNet ICCV’21 (Zhou et al., 2021) $32.70\%$
-FPN-Fusion ICRA’22 (Tomy et al., 2022) $32.28\%$
-EFNet ECCV’22 (Sun et al., 2022a) $35.33\%$
-RENet ICRA’23 (Zhou et al., 2023) $38.38\%$
-FlexEvent Ours - $\mathbf{48.64\%}$
-Figure 5: Qualitative results of state-of-the-art event camera detectors.
-We compare FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det.
-Best viewed in colors.
-See the Appendix for more examples.
-Table 4: Ablation study of components in FlexEvent.
-FlexFuser denotes the adaptive event-frame fusion module (cf.
-Sec.3.2).
-FAL denotes the frequency-adaptive learning mechanism (cf.
-Sec.
-3.3).
-The reported are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
-The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FAL.
-Modality FAL FlexFuser Frequency (Hz) Average
-$\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
-E $53.2\%$ $54.0\%$ $53.5\%$ $52.0\%$ $49.4\%$ $45.9\%$ $38.8\%$ $22.9\%$ $46.2\%$
-$54.6\%$ $54.9\%$ $54.9\%$ $54.3\%$ $53.3\%$ $50.7\%$ $44.6\%$ $30.4\%$ $49.7\%$
-E + F $\blacklozenge$ $54.9\%$ $57.3\%$ $57.7\%$ $57.8\%$ $57.2\%$ $56.1\%$ $53.7\%$ $48.3\%$ $55.4\%$
-$\mathbf{58.0\%}$ $59.6\%$ $\mathbf{60.0\%}$ $59.6\%$ $59.0\%$ $57.6\%$ $54.8\%$ $49.2\%$ $57.2\%$
-$57.4\%$ $\mathbf{60.0\%}$ $\mathbf{60.0\%}$ $\mathbf{60.1\%}$ $\mathbf{59.5\%}$ $\mathbf{58.8\%}$ $\mathbf{56.5\%}$ $\mathbf{50.9\%}$ $\mathbf{57.9\%}$
-Comparison to Multimodal Models.
-We compare FlexEvent with multimodal event-camera object detection methods such as DAGr (Gehrig & Scaramuzza, 2024), CAFR (Cao et al., 2024), and RENet (Zhou et al., 2023), which fuse event data with other sensor inputs, such as RGB frames or depth to improve detection accuracy.
+Comparisons with Multi-Modal Detectors.
+We compare FlexEvent with multimodal event-camera object detection methods such as DAGr [16], HDI-Former [31], CAFR [3], and RENet [66], which fuse event data with other sensor inputs to improve detection accuracy.
+The results are shown in Tab.
+1, Tab.
+2 and Tab.
+6.
 While these methods enhance performance over event-only approaches, they struggle to adapt to varying operational frequencies and often exhibit inadequate feature fusion in dynamic environments.
 Our approach addresses these limitations by dynamically balancing the contributions of event and frame data.
-As a result, we achieve superior performance, such as a $47.4\%$ mAP on DSEC-Detection as shown in Tab.
-2, outperforming CAFR ($38.0\%$) and RENet ($29.4\%$).
-This flexible combination of event and frame data, along with its ability to generalize across different temporal resolutions, enables our method to excel in high-frequency detection scenarios, surpassing state-of-the-art methods.
-Comparisons Across Different Categories.
-We evaluate the performance of FlexEvent across various object categories, including cars, pedestrians, and large vehicles.
+This flexible combination, along with the ability to generalize across different temporal resolutions, enables our method to excel in high-frequency event-based detection scenarios, surpassing state-of-the-art approaches.
+Generalization on High-Frequency Data.
+A key contribution of FlexEvent is its generalization capability across various operational frequencies, particularly in high-frequency scenarios.
+We evaluate this by testing detection performance at different temporal offsets, $\frac{i}{n}\Delta T$, where $n=10$, i=0,,10i=0,...,10, and $\Delta T=50$ms.
+Ground truth labels are generated by linearly interpolating object positions between frames following DAGr [16].
+In this setting, event-based methods are tested across multiple time durations, while event-frame fusion methods process one RGB frame followed by event data of varying time durations.
+As shown in Fig.
+6, existing methods degrade significantly at higher frequencies due to fixed temporal intervals and limited ability to capture fast scene changes, while our method achieves consistent, strong performance.
+Comparisons on Efficiency.
+In Tab.
+5, we compare inference times and parameter counts using an NVIDIA A50005000 2424GB GPU and an AMD EPYC $32$-Core Processor.
+Although FlexEvent comprises slightly more parameters overall, its speed matches SAST [39] and clearly surpasses DAGr [16] at all tested frequencies.
+As FlexTune is performed off-line, the high-frequency generalizability is already propagated in the model and introduces no runtime overhead.
+Our lightweight fusion head contributes only $\sim\!1.5$M parameters, while the event–frame backbone adds a further $\sim\!44$M that is already highly optimized and incurs virtually no extra latency.
+This confirms our efficiency and real-time suitability.
+Qualitative Assessments.
+We provide visual comparisons of FlexEvent and other state-of-the-art methods under different event operation frequencies, as shown in Fig.
+5.
+Unlike RVT [18] and DAGr [16], which often miss or misinterpret critical objects, our model consistently detects objects with high accuracy, even in challenging cases involving fast-moving vehicles and occluded pedestrians.
+This robustness is also evident in examples of Fig.
+7, where our method detects a pedestrian missed by RVT [18] due to sparse event data at high frequencies.
+For more detailed analyses, visualizations, and failure cases, please refer to Appendix Sec.
+C.
+Figure 7: Qualitative assessment of FlexEvent and RVT [18] under different operation frequencies.
+Figure 8: Ablation on event-frame fusion strategies under different frequencies on DSEC-Det [16].
+Table 6: Ablation study on hyperparameters.
+$\mathbf{\tau}^{\mathrm{c}}$, $\mathbf{\tau}^{\mathrm{p}}$ denotes the confidence threshold for Car and Pedestrian, respectively.
+$\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length.
+The reported results are the mAP scores on the test set of DSEC-Det [16].
+ $\mathbf{\tau}^{\mathrm{c}}$  $\mathbf{\tau}^{\mathrm{p}}$ $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) Avg
+$\mathbf{20.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
+$0.6$ $0.3$ $10$ $0.8$ $56.5$ $57.2$ $57.1$ $56.7$ $54.5$ $49.2$ $55.2$
+$0.6$ $0.3$ $10$ $0.6$ $56.7$ $57.9$ $57.7$ $57.0$ $54.3$ $47.0$ $55.1$
+$0.6$ $0.3$ 88 $0.6$ $56.3$ $59.1$ $58.8$ $58.4$ $56.2$ $\mathbf{51.2}$ $56.7$
+$0.6$ $0.3$ 66 $0.6$ $57.3$ $59.9$ $59.3$ $58.5$ $55.7$ $48.8$ $56.6$
+$0.6$ $0.6$ 66 $0.6$ $\mathbf{57.4}$ $\mathbf{60.1}$ $\mathbf{59.5}$ $\mathbf{58.8}$ $\mathbf{56.5}$ $50.9$ $\mathbf{57.2}$
+$0.8$ $0.8$ 66 $0.6$ $56.6$ $58.9$ $58.4$ $57.4$ $55.6$ $50.2$ $56.2$
+4.3 Ablation Study
+Analysis of FlexFuse.
+We assess the effect of integrating frame data via FlexFuse.
 As shown in Tab.
-2, we consistently outperform other methods, achieving $59.3\%$ mAP for cars, compared to $49.9\%$ for CAFR (Cao et al., 2024).
-This highlights its effectiveness in detecting larger, fast-moving objects, while also surpassing CAFR in pedestrian categories.
-Existing Event-based detectors struggle with detecting smaller and stationary objects due to a lack of pixel intensity changes, and fusion methods over-rely on frame data, which lacks temporal resolution.
-Our approach addresses these issues with adaptive fusion and temporal refinement, ensuring accurate detection across different object types and motion dynamics.
-This versatility reinforces its superiority over state-of-the-art methods.
-Generalization on High-Frequency Data.
-A key contribution of FlexEvent is its ability to generalize across various operational frequencies, particularly in high-frequency scenarios.
-We evaluate this by testing detection performance at different temporal offsets, $\frac{i}{n}\Delta T$, where $n=10$, $i=0,...,10$, and $\Delta T=50$ ms.
-Ground truth labels are generated by linearly interpolating object positions between frames for consistent evaluation.
-In this setting, event-based methods are tested across multiple time durations, while event-frame fusion methods process one RGB frame followed by event data of varying time durations.
-The comparison results are shown in  Fig.
-4.
-Most existing methods, such as RVT and SAST, struggle at higher frequencies due to their fixed temporal intervals and limited ability to capture fast scene changes.
-In contrast, our approach achieves $56.5\%$ mAP at $90$ Hz and $50.9\%$ at $180$ Hz.
-This improvement demonstrates that our method excels in dynamic, rapidly changing environments where accurate detection is critical for safety and reliability.
-Qualitative Assessment.
-We provide qualitative comparisons between FlexEvent and other state-of-the-art methods under different event operation frequencies, as shown in Fig.
-5 and Fig.
-6.
-Unlike RVT and DAGr, which miss critical object details, our model consistently detects objects with high accuracy, even in challenging cases involving fast-moving vehicles and occluded pedestrians.
-For instance, in Scene 2 of Fig.
-5, RVT fails to detect a pedestrian due to insufficient event data, while our approach successfully identifies the pedestrian by leveraging both frames and high-frequency event data.
-Similarly, in Scene 4, DAGr struggles with the rapid motion of a large vehicle, leading to inaccurate predictions, whereas our approach ensures precise object localization.
-In the upper instance of Fig.
-6, RVT fails to detect a pedestrian at high frequency due to insufficient event data.
-In contrast, FlexEvent successfully identifies the pedestrian by leveraging the frame-based information.
-These qualitative findings confirm that our model excels not only in quantitative metrics but also in real-world performance.
-Figure 6: Qualitative comparisons of FlexEvent and RVT (Gehrig & Scaramuzza, 2023) under different event operation frequencies.
-Our approach demonstrates a strong robustness under both low- and high-frequency scenarios.
-Best viewed in colors.
-See the Appendix for more examples.
-Table 5: Ablation study of hyperparameter configurations in the FlexEvent frameworks.
-$\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively.
-$\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length.
-The reported results are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
- $\mathbf{\tau}^{\text{car}}$  $\mathbf{\tau}^{\text{ped}}$ $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) Average
-$\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30.0}$ $\mathbf{36.0}$ $\mathbf{45.0}$ $\mathbf{60.0}$ $\mathbf{90.0}$ $\mathbf{180}$
-$0.6$ $0.3$ $10$ $0.8$ $56.5\%$ $55.9\%$ $56.7\%$ $57.2\%$ $57.1\%$ $56.7\%$ $54.5\%$ $49.2\%$ $55.5\%$
-$0.6$ $0.3$ $10$ $0.6$ $56.7\%$ $57.2\%$ $57.7\%$ $57.9\%$ $57.7\%$ $57.0\%$ $54.3\%$ $47.0\%$ $55.7\%$
-$0.6$ $0.3$ $8$ $0.6$ $56.3\%$ $58.5\%$ $58.8\%$ $59.1\%$ $58.8\%$ $58.4\%$ $56.2\%$ $\mathbf{51.2\%}$ $57.2\%$
-$0.6$ $0.3$ $6$ $0.6$ $57.3\%$ $59.4\%$ $59.7\%$ $59.9\%$ $59.3\%$ $58.5\%$ $55.7\%$ $48.8\%$ $57.3\%$
-$0.6$ $0.6$ $6$ $0.6$ $\mathbf{57.4\%}$ $\mathbf{60.0}\%$ $\mathbf{60.0\%}$ $\mathbf{60.1\%}$ $\mathbf{59.5\%}$ $\mathbf{58.8\%}$ $\mathbf{56.5\%}$ $50.9\%$ $\mathbf{57.9\%}$
-$0.8$ $0.8$ $6$ $0.6$ $56.6\%$ $58.7\%$ $59.1\%$ $58.9\%$ $58.4\%$ $57.4\%$ $55.6\%$ $50.2\%$ $56.9\%$
-4.3 Ablation studies
-Component Analysis.
-We conduct ablation studies by selectively removing key modules: the FlexFuser and FAL mechanisms.
+5, adding frame information boosts average mAP from $43.7\%$ to $56.4\%$, with greater gains at higher frequencies where event data alone becomes sparse.
+Our adaptive gating fuses event and frame features more effectively than simple Add, Concat, or vanilla Attention (Fig.
+8), achieving the best accuracy across all frequencies and providing robust representations under diverse conditions.
+Analysis of FlexTune.
+In parallel, we assess FlexTune, focusing on scenarios where event data are sparse and fast-evolving.
+Without it, performance at high frequencies collapses, as seen in the first row of Tab.
+5 where mAP at $180$ Hz is only $22.9\%$.
+However, activating FlexTune raises this score to $30.4\%$, demonstrating the value of iterative self-training and refinement for frequency adaptation.
+We also test training with interpolated high-frequency labels, but this approach struggles with objects that appear or disappear rapidly, reducing recall of the label.
+By contrast, FlexTune updates the model with refined, temporally consistent labels, improving detection quality.
+Hyperparameter Tuning.
+We investigate key hyperparameters in the FlexTune mechanism, specifically, the confidence thresholds for cars $\mathbf{\tau}^{\mathrm{c}}$ and pedestrians $\mathbf{\tau}^{\mathrm{p}}$, the IoU threshold for bounding-box association $\mathbf{\tau}^{\text{iou}}$, and the minimum track length for temporal refinement $\mathbf{L}^{\text{track}}$.
 As shown in Tab.
-4, removing the FAL mechanism results in a significant performance drop, particularly in high-frequency scenarios, underscoring its role in adapting to varying frequencies and generating frequency-adjusted labels.
-Similarly, omitting the FlexFuser module leads to a marked decrease in mAP, highlighting the importance of adaptive event-frame fusion for accurate detection across different operational frequencies.
-We also test training with interpolated labels for high-frequency testing, but this approach reduces recall by missing objects that suddenly appear or disappear, making it less effective than FAL.
-Hyperparameter Tuning.
-We tune the hyperparameters of the FAL mechanism, focusing on key settings like the confidence threshold ($\tau$), IoU threshold, and track length for temporal refinement.
-As shown in Tab.
-5, lowering the confidence threshold improves recall but reduces precision, as the model becomes more lenient in detecting objects.
-Applying overly strict conditions, such as a higher confidence threshold or IoU threshold, lowers recall by filtering out valid detections.
-The optimal configuration is achieved with $\tau=0.6$ and a track length of 6, balancing precision and recall for both low- and high-frequency conditions.
-These moderate settings ensure that FlexEvent maintains robust performance and stable detection accuracy across diverse environments.
+8, lowering confidence thresholds can improve recall by admitting more detections, but it risks increasing false positives and thus reducing overall precision.
+Conversely, setting overly strict thresholds, such as a higher IoU requirement or confidence level, might filter out potential detections, lowering recall.
+An intermediate setting of $\tau=0.6$ for both classes, with a track length of $6$ and an IoU threshold of $0.6$, emerges as the best trade-off, balancing recall and precision across both low- and high-frequency conditions.
+These moderate, carefully tuned parameters ensure that FlexEvent remains robust and accurate in a range of real-world scenarios.
 5 Conclusion
-This paper introduces FlexEvent, an event-camera object detection framework designed to operate across arbitrary frequencies.
-By combining FlexFuser for adaptive event-frame fusion and FAL for frequency-adaptive learning, we leverage the rich temporal information of event data and the semantic detail of RGB frames to overcome the limitations of existing methods, providing a flexible solution for dynamic environments.
-Extensive experiments on large-scale datasets show that our approach significantly outperforms state-of-the-art methods, particularly in high-frequency scenarios, demonstrating its robustness and adaptability for real-world applications such as autonomous driving and robotics.
+This paper introduces FlexEvent, an event-camera detection framework designed to operate across varying frequencies.
+By combining FlexFuse for adaptive event-frame fusion and FlexTune for frequency-adaptive learning, we leverage the rich temporal information of event data and the semantic detail of RGB frames to overcome the limitations of existing methods, providing a flexible solution for dynamic environments.
+Extensive experiments on large-scale datasets show that our approach significantly outperforms state-of-the-art methods, particularly in high-frequency scenarios, demonstrating its robustness and adaptability for real-world applications.
 Appendix
-In this appendix, we supplement the following materials to support the findings and conclusions drawn in the main body of this paper.
-[…]
-[…]
-Appendix A A Additional Implementation Details
+Appendix A Additional Implementation Details
 In this section, to facilitate future reproductions, we elaborate on the necessary details in terms of the datasets, evaluation metrics, and implementation details adopted in our experiments.
 A.1 Datasets
-In this work, we develop and validate our proposed approach on the large-scale DSEC dataset (Gehrig et al., 2021b).
+In this work, we develop and validate our proposed approach on the large-scale DSEC dataset [17].
 DSEC serves as a high-resolution, large-scale multimodal dataset designed to capture real-world driving scenarios under various conditions.
@@
 This alignment ensures that the event data and RGB frames are temporally and spatially synchronized.
-Table 6: Summary of key statistics from the event camera object detection datasets used in this work.
-Dataset Reference Classes Frames Sequences Class Names
-DSEC-MOD (Zhou et al., 2023) $1$ $13,314$ $16$ Car
-DSEC-Detection (Tomy et al., 2022) $3$ $52,727$ $41$ Car
+In our experiments, we utilize three comprehensive versions of DSEC tailored for object detection: DSEC-Det [16], DSEC-Detection [50], and DSEC-MOD [66].
+A summary of the key statistics of these datasets is listed in Tab.
+7.
+  • DSEC-Det [16]: This version is developed by the original DSEC team and includes annotations generated using the QDTrack multi-object tracker [11, 38].
+The annotation process combines automated multi-object tracking with manual refinement to ensure high-quality and accurate detection labels.
+Compared to the original DSEC dataset, DSEC-Det [16] introduces additional sequences specifically designed to capture complex and dynamic urban environments, featuring crowded pedestrian areas, moving vehicles, and diverse lighting conditions [29].
+These challenging scenarios provide a rich testing ground for evaluating object detection algorithms in real-world driving settings.
+DSEC-Det [16] comprises $60$ sequences and 78,34478{,}344 frames, making it the most extensive dataset used in this study.
+It captures diverse urban scenes with dynamic elements, such as crowded pedestrian areas and moving vehicles.
+Covering eight object categories relevant to autonomous driving, Car, Pedestrian, Bus, Bicycle, Truck, Motorcycle, Rider, and Train, the dataset provides a robust foundation for training and evaluating object detection models across various driving scenarios.
+In our experiments on DSEC-Det [16], to maintain consistency with the experimental setup of previous work DAGr [16], we report results on two categories: Car and Pedestrian.
+  • DSEC-Detection [50]: This dataset comprises $41$ sequences with a total of 52,72752{,}727 frames.
+Focusing on three fundamental object categories – Car, Pedestrian, and Large Vehicle – this version emphasizes high-precision annotations for these critical classes in autonomous driving.
+The initial annotations are generated using the YOLOv5 model [26] on RGB frames, leveraging its robust performance in real-time object detection.
+These annotations are then transferred to the corresponding event frames through homographic transformation, ensuring spatial alignment between the two modalities.
+A subsequent manual refinement process corrects any discrepancies and enhances annotation quality, resulting in a dataset that provides accurate and reliable labels for event-based object detection.
+  • DSEC-MOD [66]: This dataset extends the object detection capabilities to moving-object detection across diverse urban environments.
+It includes $16$ sequences containing 13,31413{,}314 frames and is specifically focused on the Car category, making it highly suitable for complex detection tasks in varied urban settings, such as intersections, highways, and residential areas.
+Featuring high-frequency and dense annotations, the dataset provides a valuable resource for evaluating the performance of event-based object detectors under challenging real-world conditions.
+These three versions of the DSEC dataset together provide a comprehensive platform for benchmarking and evaluating event-based object detection methods, capturing a wide spectrum of scenarios, object categories, and environmental conditions.
+Among them, DSEC-Det [16] is the largest, most recent, and most comprehensive, annotated, and released by the original DSEC authors.
+Thus, we prioritize it as the primary benchmark for reporting results, ensuring relevance and reliability.
+DSEC-Detection [50] and DSEC-MOD [66] are datasets used by two recent event-frame fusion methods, CAFR [3] and RENet [66].
+To validate the effectiveness of our method, we also report results on these two datasets.
+Table 7: Summary of key statistics from the three event datasets [16, 50, 66] used in this work.
+Dataset Classes Frames Sequences Class Names
+DSEC-MOD [66] 11 $13,314$ $16$ Car
+DSEC-Detection [50] 33 $52,727$ $41$ Car
 Pedestrian
 Large-Vehicle
-DSEC-Det (Gehrig & Scaramuzza, 2024) $8$ $78,344$ $60$ Car
+DSEC-Det [16] 88 $78,344$ $60$ Car
 Pedestrian
@@
 Train
-In our experiments, we utilize three comprehensive versions of DSEC tailored for object detection: DSEC-Det (Gehrig & Scaramuzza, 2024), DSEC-Detection (Tomy et al., 2022), and DSEC-MOD (Zhou et al., 2023).
-A summary of the key statistics of these datasets is listed in Tab.
-6.
-  • DSEC-Det (Gehrig & Scaramuzza, 2024): This version is developed by the original DSEC team and includes annotations generated using the QDTrack multi-object tracker (Fischer et al., 2023; Pang et al., 2021).
-The annotation process combines automated multi-object tracking with manual refinement to ensure high-quality and accurate detection labels.
-Compared to the original DSEC dataset, DSEC-Det introduces additional sequences specifically designed to capture complex and dynamic urban environments, featuring crowded pedestrian areas, moving vehicles, and diverse lighting conditions.
-These challenging scenarios provide a rich testing ground for evaluating object detection algorithms in real-world driving settings.
-DSEC-Det comprises $60$ sequences and $78,344$ frames, making it the most extensive dataset used in this study.
-It captures diverse urban scenes with dynamic elements, such as crowded pedestrian areas and moving vehicles.
-Covering eight object categories relevant to autonomous driving – Car, Pedestrian, Bus, Bicycle, Truck, Motorcycle, Rider, and Train – the dataset provides a robust foundation for training and evaluating object detection models across various driving scenarios.
-In our experiments on DSEC-Det, to maintain consistency with the experimental setup of previous work DAGr (Gehrig & Scaramuzza, 2024), we report results on two categories: Car and Pedestrian.
-  • DSEC-Detection (Tomy et al., 2022): This dataset comprises $41$ sequences with a total of $52,727$ frames.
-Focusing on three fundamental object categories – Car, Pedestrian, and Large Vehicle – this version emphasizes high-precision annotations for these critical classes in autonomous driving.
-The initial annotations are generated using the YOLOv5 model (Jocher, 2020) on RGB frames, leveraging its robust performance in real-time object detection.
-These annotations are then transferred to the corresponding event frames through homographic transformation, ensuring spatial alignment between the two modalities.
-A subsequent manual refinement process corrects any discrepancies and enhances annotation quality, resulting in a dataset that provides accurate and reliable labels for event-based object detection.
-  • DSEC-MOD (Zhou et al., 2023): This dataset extends the object detection capabilities to moving-object detection across diverse urban environments.
-It includes $16$ sequences containing $13,314$ frames and is specifically focused on the Car category, making it highly suitable for complex detection tasks in varied urban settings, such as intersections, highways, and residential areas.
-Featuring high-frequency and dense annotations, the dataset provides a valuable resource for evaluating the performance of event-based object detectors under challenging real-world conditions.
-These three versions of the DSEC dataset together provide a comprehensive platform for benchmarking and evaluating event-based object detection methods, capturing a wide spectrum of scenarios, object categories, and environmental conditions.
-Among them, DSEC-Det is the largest, most recent and most comprehensive, annotated and released by the original DSEC authors.
-Thus, we prioritize it as the primary benchmark for reporting results, ensuring relevance and reliability.
-DSEC-Detection and DSEC-MOD are datasets used by two recent event-frame fusion methods, CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023).
-To validate the effectiveness of our method, we also report results on these two datasets.
 A.2 Baselines
@@
   • Event-Only Methods.
-We include state-of-the-art event-only object detectors, namely RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), LEOD (Wu et al., 2024), and SSM (Zubić et al., 2024), which are originally trained on event-only datasets like Gen1 (De Tournemire et al., 2020) and 1Mpx (Perot et al., 2020).
-To ensure a fair comparison, we retrain these methods on the DSEC-Det dataset following their respective training protocols.
+We include state-of-the-art event-only object detectors, namely RVT [18], SAST [39], LEOD [55], and SSM [70], which are originally trained on event-only datasets like Gen11 [10] and 11Mpx [40].
+To ensure a fair comparison, we retrain these methods on the DSEC-Det [16] dataset following their respective training protocols.
   • Event-Frame Fusion Methods.
-For event-frame fusion methods on DSEC-Det, we include DAGr, as it has been evaluated on this dataset.
-We report the scores of DAGr (Gehrig & Scaramuzza, 2024) from the original paper to ensure consistency and fairness.
-For the DSEC-Detection and DSEC-MOD datasets, we train our model following the standard training and evaluation settings, and compare it against state-of-the-art methods CAFR (Cao et al., 2024) and RENet (Zhou et al., 2023), as reported in their respective papers.
-For other methods evaluated on DSEC-Detection and DSEC-MOD, we reference the results reported in the CAFR and RENet papers, respectively.
-Since DAGr’s training code is not publicly available, we could not reproduce its results on DSEC-Detection and DSEC-MOD.
-These comparisons ensure a fair and comprehensive evaluation while adhering to resource and code availability constraints.
+For event-frame fusion methods on DSEC-Det [16], we include DAGr [16] and HDI-Former [31], as they have been evaluated on this dataset.
+We report the scores of DAGr and HDI-Former from the original paper to ensure consistency and fairness.
+For the DSEC-Detection [50] and DSEC-MOD [66] datasets, we train our model following the standard training and evaluation settings, and compare it against state-of-the-art methods CAFR [3] and RENet [66], as reported in their respective papers.
+For other methods evaluated on DSEC-Detection [50] and DSEC-MOD [66], we reference the results reported in the CAFR and RENet papers, respectively.
+These comparisons among state-of-the-art works ensure a fair and comprehensive evaluation while adhering to resource and code availability constraints.
 A.3 Evaluation Metrics
@@
 Mathematically, the Average Precision (AP) for a single class is calculated as:
-[…]
+$\text{AP}=\int_{0}^{1}p(r)\,dr~,$ (10)
 where $p(r)$ represents the precision at a given recall level $r$.
@@
 This provides an overall measure of model performance across different levels of localization precision.
-In addition to mAP, we also report the following metrics from the COCO evaluation protocol (Lin et al., 2014):
+In addition to mAP, we also report the following metrics from the COCO evaluation protocol [35]:
   • AP50: The average precision when evaluated at a fixed IoU threshold of $0.50$, indicating how well the model performs with relatively lenient localization criteria.
@@
 By reporting these metrics, we obtain a more nuanced understanding of the model’s detection capabilities across varying object sizes and localization precision levels, ensuring a comprehensive evaluation of detection performance.
-A.4 Training & Inference Details
-We train our models using mixed precision to optimize both memory efficiency and training speed.
-The training process spans 100,000 iterations, utilizing the ADAM optimizer (Kingma, 2014) with a OneCycle learning rate schedule (Smith & Topin, 2019), which gradually decays from a peak learning rate to enhance convergence.
-Consistent with RVT (Gehrig & Scaramuzza, 2023), we employ a mixed batching strategy to balance computational efficiency and memory usage. Specifically:
-  • Standard Backpropagation Through Time (BPTT): Applied to half of the training samples, allowing for full sequence training.
-  • Truncated BPTT (TBPTT): Used for the other half, reducing memory usage by splitting sequences into smaller segments.
-For data augmentation, we apply random horizontal flipping and zoom transformations (both zoom-in and zoom-out) to enhance the diversity of training samples.
-Our training process utilize the YOLOX framework (Zheng et al., 2021), a versatile object detection framework known for its efficient and high-performing architecture.
-We employ a multi-component loss function to optimize our model effectively:
-  • Intersection over Union (IoU) Loss: This loss component measures the overlap between the predicted bounding boxes and the ground-truth boxes, ensuring that the predicted regions closely match the actual object locations.
-  • Classification Loss: This component evaluates the accuracy of class predictions for each detected object, ensuring that the model correctly identifies the category of each detected instance.
-  • Regression Loss: This loss assesses the precision of the predicted bounding box coordinates, helping the model refine the location and size of bounding boxes to align closely with the ground-truth annotations.
-To ensure stable training, these loss components are averaged across both the batch and sequence length at each optimization step.
-This averaging process helps to reduce variance during training and facilitates smoother convergence of the model parameters.
-Training Configuration.
-The training is conducted with a batch size of 8, which provides an optimal balance between efficient GPU utilization and memory requirements.
-Each training sample contains a sequence length of 11 frames, allowing the model to learn temporal dependencies effectively.
-The frame backbone’s weights are initialized using pre-trained ResNet and the event backbone’s weights are initialized using pre-trained RVT.
-The learning rate is set to $1\times 10^{-4}$, following a OneCycle learning rate schedule that allows efficient exploration of the learning space and helps in achieving faster convergence.
-Hardware & Training Time.
-All training experiments are carried out on two NVIDIA RTX A5000 GPUs, each with 24GB of memory, providing the computational resources necessary for handling the high-resolution event data and RGB frames.
-The complete training process, including all iterations and model optimization, takes approximately one day, demonstrating the efficiency of our implementation in terms of both training speed and resource utilization.
-Table 7: Comparative efficiency analysis of state-of-the-art event camera detectors on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024), comparing both event-only and event-frame fusion methods.
+Table 8: The complete results of the efficiency analysis of state-of-the-art event detectors on the test set of DSEC-Det [16], comparing both event-only and event-frame fusion methods.
 This table reports inference times at various frequencies, measured in milliseconds (ms).
-Modality Method Param Frequency (Hz)
-(M) $\mathbf{20.0}$ $\mathbf{27.5}$ $\mathbf{30}$ $\mathbf{36}$ $\mathbf{45}$ $\mathbf{60}$ $\mathbf{90}$ $\mathbf{180}$ $\mathbf{200}$
-E RVT $18.5$ $9.20$ $8.67$ $8.35$ $7.93$ $7.61$ $7.51$ $7.19$ $6.77$ $6.34$
-SAST 18.9 $14.06$ $13.11$ $12.68$ $12.37$ $11.95$ $11.63$ $11.52$ $11.10$ $10.36$
-SSM 18.2 $8.79$ $8.26$ $8.08$ $7.71$ $7.55$ $7.30$ $6.90$ $6.54$ $6.12$
-E + F DAGr-50 $34.6$ $73.35$ $65.73$ $60.02$ $55.11$ $51.00$ $48.00$ $45.29$ $43.89$ $37.58$
-FlexEvent $45.4$ $14.27$ $13.53$ $13.32$ $13.00$ $12.79$ $12.58$ $12.47$ $12.37$ $12.12$
-Table 8: The complete results of the ablation study (cf.
-Tab.
-4) of different components in the FlexEvent framework.
-FlexFuser denotes the adaptive event-frame fusion module (cf.
-Sec.3.2).
-FAL denotes the frequency-adaptive learning module (cf.
-Sec.
-3.3).
-The reported results are the mAP, AP50, AP75, APS, APM, and APL scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
-The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FAL.
-The best and 2nd best scores of each metric are highlighted in bold and underline, respectively.
-Modality FAL FlexFuser Frequency (Hz) mAP AP50 AP75 APS APM APL
+Modal Method Param Operational Frequency
+(M) $\mathbf{20.0}$ Hz $\mathbf{36}$ Hz $\mathbf{45}$ Hz $\mathbf{60}$ Hz $\mathbf{90}$ Hz $\mathbf{180}$ Hz $\mathbf{200}$ Hz
+E RVT [18] $18.5$ $9.20$ ms $7.93$ ms $7.61$ ms $7.51$ ms $7.19$ ms $6.77$ ms $6.34$ ms
+SAST [39] $18.9$ $14.06$ ms $12.37$ ms $11.95$ ms $11.63$ ms $11.52$ ms $11.10$ ms $10.36$ ms
+SSM [70] $18.2$ $8.79$ ms $7.71$ ms $7.55$ ms $7.30$ ms $6.90$ ms $6.54$ ms $6.12$ ms
+E + F DAGr-50 [16] $34.6$ $73.35$ ms $55.11$ ms $51.00$ ms $48.00$ ms $45.29$ ms $43.89$ ms $37.58$ ms
+FlexEvent $45.4$ $14.27$ ms $13.00$ ms $12.79$ ms $12.58$ ms $12.47$ ms $12.37$ ms $12.12$ ms
+Table 9: The complete ablation results of different components in the FlexEvent framework.
+FlexFuse denotes the adaptive event-frame fusion module.
+FlexTune denotes the FlexTune learning mechanism.
+The reported results are the mAP, AP50, AP75, APS, APM, and APL scores on the test set of DSEC-Det [16].
+The symbol $\blacklozenge$ denotes the use of interpolated ground truth labels at high frequencies in FlexTune.
+The best and 2nd best scores of each metric are highlighted in bold and underline.
+Modality FlexTune FlexFuse Frequency (Hz) mAP AP50 AP75 APS APM APL
 Event $20.0$ $53.2\%$ $\mathbf{77.2\%}$ $58.1\%$ $\mathbf{46.4\%}$ $64.4\%$ $83.0\%$
@@
 $36.0$ $54.3\%$ $77.1\%$ $60.5\%$ $46.8\%$ $66.7\%$ $83.4\%$
-$45.0$ $53.3\%$ $75.3\%$ $59.8\%$ $45.6\%$ $65.4\%$ $\mathbf{83.8\%}$
+$45.0$ $53.3\%$ $75.3\%$ $59.8\%$ $\% 65.465.4\%$ $\mathbf{83.8\%}$
 $60.0$ $50.7\%$ $72.4\%$ $57.3\%$ $42.3\%$ $63.5\%$ $83.5\%$
@@
 $\blacklozenge$ $36.0$ $\mathbf{57.8\%}$ $75.7\%$ $66.5\%$ $52.5\%$ $67.9\%$ $87.2\%$
-$\blacklozenge$ $45.0$ $57.2\%$ $75.5\%$ $65.4\%$ $51.6\%$ $\mathbf{68.2\%}$ $\mathbf{87.5\%}$
+$\blacklozenge$ $45.0$ $2\% 75.57$$5\%$ $65.4\%$ $51.6\%$ $\mathbf{68.2\%}$ $\mathbf{87.5\%}$
 $\blacklozenge$ $60.0$ $56.1\%$ $74.2\%$ $63.4\%$ $50.1\%$ $68.1\%$ $86.5\%$
@@
 $180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
-Table 9: The complete results of the ablation study (cf.
-Tab.
-5) of different hyperparameter configurations in the FlexEvent framework.
+Table 10: The complete ablation results of different hyperparameter configurations in the FlexEvent framework.
 $\mathbf{\tau}^{\text{car}}$, $\mathbf{\tau}^{\text{ped}}$ denotes the confidence threshold for car and pedestrian, respectively.
 $\mathbf{\tau}^{\text{iou}}$ denotes the IoU threshold when filter by tracking, $\mathbf{L}^{\text{track}}$ denotes the minimum track length.
-The reported results are the mAP scores on the test set of DSEC-Det (Gehrig & Scaramuzza, 2024).
-The best and 2nd best scores of each metric from each hyperparameter configuration are highlighted in bold and underline, respectively.
+The reported results are the mAP scores on the test set of DSEC-Det [16].
+The best and 2nd best scores of each metric from each hyperparameter configuration are highlighted in bold and underline.
  $\mathbf{\tau}^{\text{car}}$  $\mathbf{\tau}^{\text{ped}}$  $\mathbf{L}^{\text{track}}$  $\mathbf{\tau}^{\text{iou}}$ Frequency (Hz) mAP AP50 AP75 APS APM APL
@@
 $0.6$ $0.3$ $10$ $0.6$ $180.0$ $47.0\%$ $69.0\%$ $49.1\%$ $37.8\%$ $61.1\%$ $83.9\%$
-$0.6$ $0.3$ $8$ $0.6$ $20.0$ $56.3\%$ $77.2\%$ $64.9\%$ $50.4\%$ $64.1\%$ $83.3\%$
-$0.6$ $0.3$ $8$ $0.6$ $27.5$ $58.5\%$ $78.3\%$ $68.1\%$ $52.5\%$ $66.8\%$ $84.8\%$
-$0.6$ $0.3$ $8$ $0.6$ $30.0$ $58.8\%$ $78.5\%$ $68.7\%$ $52.6\%$ $67.9\%$ $85.7\%$
-$0.6$ $0.3$ $8$ $0.6$ $36.0$ $\mathbf{59.1\%}$ $\mathbf{78.8\%}$ $\mathbf{69.0\%}$ $\mathbf{52.7\%}$ $68.8\%$ $\mathbf{86.5\%}$
-$0.6$ $0.3$ $8$ $0.6$ $45.0$ $58.8\%$ $78.2\%$ $68.6\%$ $52.3\%$ $68.8\%$ $86.1\%$
-$0.6$ $0.3$ $8$ $0.6$ $60.0$ $58.4\%$ $77.9\%$ $67.5\%$ $51.3\%$ $\mathbf{69.6\%}$ $85.5\%$
-$0.6$ $0.3$ $8$ $0.6$ $90.0$ $56.2\%$ $76.6\%$ $64.8\%$ $48.5\%$ $68.3\%$ $84.7\%$
-$0.6$ $0.3$ $8$ $0.6$ $180.0$ $51.2\%$ $71.9\%$ $56.3\%$ $42.6\%$ $64.2\%$ $82.9\%$
-$0.6$ $0.3$ $6$ $0.6$ $20.0$ $57.3\%$ $80.0\%$ $65.2\%$ $51.2\%$ $65.8\%$ $84.1\%$
-$0.6$ $0.3$ $6$ $0.6$ $27.5$ $59.4\%$ $81.3\%$ $68.5\%$ $53.4\%$ $68.8\%$ $\mathbf{85.7\%}$
-$0.6$ $0.3$ $6$ $0.6$ $30.0$ $59.7\%$ $\mathbf{81.7\%}$ $\mathbf{69.0\%}$ $\mathbf{53.7\%}$ $69.0\%$ $85.3\%$
-$0.6$ $0.3$ $6$ $0.6$ $36.0$ $\mathbf{59.9\%}$ $81.4\%$ $\mathbf{69.0\%}$ $53.6\%$ $\mathbf{69.8\%}$ $\mathbf{85.7\%}$
-$0.6$ $0.3$ $6$ $0.6$ $45.0$ $59.3\%$ $80.5\%$ $67.9\%$ $52.8\%$ $69.4\%$ $85.6\%$
-$0.6$ $0.3$ $6$ $0.6$ $60.0$ $58.5\%$ $79.6\%$ $67.0\%$ $51.5\%$ $69.6\%$ $84.9\%$
-$0.6$ $0.3$ $6$ $0.6$ $90.0$ $55.7\%$ $77.2\%$ $62.8\%$ $48.1\%$ $67.9\%$ $84.3\%$
-$0.6$ $0.3$ $6$ $0.6$ $180.0$ $48.8\%$ $70.6\%$ $50.9\%$ $40.8\%$ $61.1\%$ $83.4\%$
-$0.6$ $0.6$ $6$ $0.6$ $20.0$ $57.4\%$ $78.2\%$ $66.6\%$ $51.7\%$ $64.9\%$ $83.7\%$
-$0.6$ $0.6$ $6$ $0.6$ $27.5$ $60.0\%$ $79.4\%$ $70.1\%$ $53.5\%$ $68.4\%$ $\mathbf{86.1\%}$
-$0.6$ $0.6$ $6$ $0.6$ $30.0$ $60.0\%$ $\mathbf{79.7\%}$ $\mathbf{70.8\%}$ $\mathbf{53.6\%}$ $69.9\%$ $\mathbf{86.1\%}$
-$0.6$ $0.6$ $6$ $0.6$ $36.0$ $\mathbf{60.1\%}$ $79.6\%$ $\mathbf{70.8\%}$ $53.2\%$ $70.3\%$ $85.7\%$
-$0.6$ $0.6$ $6$ $0.6$ $45.0$ $59.5\%$ $79.0\%$ $69.5\%$ $52.5\%$ $70.8\%$ $85.3\%$
-$0.6$ $0.6$ $6$ $0.6$ $60.0$ $58.8\%$ $78.5\%$ $69.0\%$ $51.1\%$ $\mathbf{71.1\%}$ $85.3\%$
-$0.6$ $0.6$ $6$ $0.6$ $90.0$ $56.5\%$ $76.5\%$ $65.4\%$ $48.2\%$ $70.1\%$ $83.8\%$
-$0.6$ $0.6$ $6$ $0.6$ $180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
-$0.8$ $0.8$ $6$ $0.6$ $20.0$ $56.6\%$ $80.7\%$ $65.5\%$ $50.8\%$ $65.4\%$ $82.6\%$
-$0.8$ $0.8$ $6$ $0.6$ $27.5$ $58.7\%$ $81.9\%$ $68.9\%$ $\mathbf{52.8\%}$ $68.6\%$ $84.9\%$
-$0.8$ $0.8$ $6$ $0.6$ $30.0$ $\mathbf{59.1\%}$ $\mathbf{82.0\%}$ $\mathbf{69.2\%}$ $52.7\%$ $69.5\%$ $84.8\%$
-$0.8$ $0.8$ $6$ $0.6$ $36.0$ $58.9\%$ $81.7\%$ $68.8\%$ $52.6\%$ $\mathbf{69.9\%}$ $85.0\%$
-$0.8$ $0.8$ $6$ $0.6$ $45.0$ $58.4\%$ $81.4\%$ $67.7\%$ $51.6\%$ $69.8\%$ $\mathbf{85.1\%}$
-$0.8$ $0.8$ $6$ $0.6$ $60.0$ $57.4\%$ $80.1\%$ $66.6\%$ $50.2\%$ $\mathbf{69.9\%}$ $84.3\%$
-$0.8$ $0.8$ $6$ $0.6$ $90.0$ $55.7\%$ $78.6\%$ $63.7\%$ $47.8\%$ $68.9\%$ $84.2\%$
-$0.8$ $0.8$ $6$ $0.6$ $180.0$ $50.2\%$ $74.0\%$ $55.2\%$ $41.5\%$ $63.7\%$ $83.1\%$
-Appendix B B Additional Quantitative Results
-B.1 Comparison of Efficiency
-We present a comparative analysis of inference times and parameter counts for the evaluated methods in Tab.
-7.
-All experiments are conducted on an NVIDIA RTX A5000 24GB GPU paired with an AMD EPYC 9354P 32-Core Processor operating at 3.8 GHz.
-The results demonstrate that, despite having a higher parameter count, FlexEvent achieves inference times comparable to the event-only method SAST and significantly outperforms the event-frame fusion method DAGr in terms of speed.
-Additionally, FlexEvent consistently surpasses DAGr in inference time across all tested frequencies.
-These results underscore the efficiency and rapid performance of FlexEvent, highlighting its suitability for real-time applications.
+$0.6$ $0.3$ 88 $0.6$ $20.0$ $56.3\%$ $77.2\%$ $64.9\%$ $50.4\%$ $64.1\%$ $83.3\%$
+$0.6$ $0.3$ 88 $0.6$ $27.5$ $58.5\%$ $78.3\%$ $68.1\%$ $52.5\%$ $66.8\%$ $84.8\%$
+$0.6$ $0.3$ 88 $0.6$ $30.0$ $58.8\%$ $78.5\%$ $68.7\%$ $52.6\%$ $67.9\%$ $85.7\%$
+$0.6$ $0.3$ 88 $0.6$ $36.0$ $\mathbf{59.1\%}$ $\mathbf{78.8\%}$ $\mathbf{69.0\%}$ $\mathbf{52.7\%}$ $68.8\%$ $\mathbf{86.5\%}$
+$0.6$ $0.3$ 88 $0.6$ $45.0$ $58.8\%$ $78.2\%$ $68.6\%$ $52.3\%$ $68.8\%$ $86.1\%$
+$0.6$ $0.3$ 88 $0.6$ $60.0$ $58.4\%$ $77.9\%$ $67.5\%$ $51.3\%$ $\mathbf{69.6\%}$ $85.5\%$
+$0.6$ $0.3$ 88 $0.6$ $90.0$ $56.2\%$ $76.6\%$ $64.8\%$ $48.5\%$ $68.3\%$ $84.7\%$
+$0.6$ $0.3$ 88 $0.6$ $180.0$ $51.2\%$ $71.9\%$ $56.3\%$ $\% 64.264.2\%$ $82.9\%$
+$0.6$ $0.3$ 66 $0.6$ $20.0$ $57.3\%$ $80.0\%$ $65.2\%$ $51.2\%$ $65.8\%$ $84.1\%$
+$0.6$ $0.3$ 66 $0.6$ $27.5$ $59.4\%$ $81.3\%$ $68.5\%$ $53.4\%$ $68.8\%$ $\mathbf{85.7\%}$
+$0.6$ $0.3$ 66 $0.6$ $30.0$ $59.7\%$ $\mathbf{81.7\%}$ $\mathbf{69.0\%}$ $\mathbf{53.7\%}$ $69.0\%$ $85.3\%$
+$0.6$ $0.3$ 66 $0.6$ $36.0$ $\mathbf{59.9\%}$ $81.4\%$ $\mathbf{69.0\%}$ $53.6\%$ $\mathbf{69.8\%}$ $\mathbf{85.7\%}$
+$0.6$ $0.3$ 66 $0.6$ $45.0$ $59.3\%$ $80.5\%$ $67.9\%$ $52.8\%$ $69.4\%$ $85.6\%$
+$0.6$ $0.3$ 66 $0.6$ $60.0$ $58.5\%$ $79.6\%$ $67.0\%$ $51.5\%$ $69.6\%$ $84.9\%$
+$0.6$ $0.3$ 66 $0.6$ $90.0$ $55.7\%$ $77.2\%$ $62.8\%$ $48.1\%$ $67.9\%$ $84.3\%$
+$0.6$ $0.3$ 66 $0.6$ $180.0$ $48.8\%$ $70.6\%$ $50.9\%$ $40.8\%$ $61.1\%$ $83.4\%$
+$0.6$ $0.6$ 66 $0.6$ $20.0$ $57.4\%$ $78.2\%$ $66.6\%$ $51.7\%$ $64.9\%$ $83.7\%$
+$0.6$ $0.6$ 66 $0.6$ $27.5$ $60.0\%$ $79.4\%$ $70.1\%$ $53.5\%$ $68.4\%$ $\mathbf{86.1\%}$
+$0.6$ $0.6$ 66 $0.6$ $30.0$ $60.0\%$ $\mathbf{79.7\%}$ $\mathbf{70.8\%}$ $\mathbf{53.6\%}$ $69.9\%$ $\mathbf{86.1\%}$
+$0.6$ $0.6$ 66 $0.6$ $36.0$ $\mathbf{60.1\%}$ $79.6\%$ $\mathbf{70.8\%}$ $53.2\%$ $70.3\%$ $85.7\%$
+$0.6$ $0.6$ 66 $0.6$ $45.0$ $59.5\%$ $79.0\%$ $69.5\%$ $52.5\%$ $70.8\%$ $85.3\%$
+$0.6$ $0.6$ 66 $0.6$ $60.0$ $58.8\%$ $78.5\%$ $69.0\%$ $51.1\%$ $\mathbf{71.1\%}$ $85.3\%$
+$0.6$ $0.6$ 66 $0.6$ $90.0$ $56.5\%$ $76.5\%$ $65.4\%$ $48.2\%$ $70.1\%$ $83.8\%$
+$0.6$ $0.6$ 66 $0.6$ $180.0$ $50.9\%$ $71.4\%$ $56.2\%$ $41.6\%$ $65.4\%$ $82.9\%$
+$0.8$ $0.8$ 66 $0.6$ $20.0$ $56.6\%$ $80.7\%$ $65.5\%$ $50.8\%$ $65.4\%$ $82.6\%$
+$0.8$ $0.8$ 66 $0.6$ $27.5$ $58.7\%$ $81.9\%$ $68.9\%$ $\mathbf{52.8\%}$ $68.6\%$ $84.9\%$
+$0.8$ $0.8$ 66 $0.6$ $30.0$ $\mathbf{59.1\%}$ $\mathbf{82.0\%}$ $\mathbf{69.2\%}$ $52.7\%$ $69.5\%$ $84.8\%$
+$0.8$ $0.8$ 66 $0.6$ $36.0$ $58.9\%$ $81.7\%$ $68.8\%$ $52.6\%$ $\mathbf{69.9\%}$ $85.0\%$
+$0.8$ $0.8$ 66 $0.6$ $45.0$ $58.4\%$ $81.4\%$ $67.7\%$ $51.6\%$ $69.8\%$ $\mathbf{85.1\%}$
+$0.8$ $0.8$ 66 $0.6$ $60.0$ $57.4\%$ $80.1\%$ $66.6\%$ $50.2\%$ $\mathbf{69.9\%}$ $84.3\%$
+$0.8$ $0.8$ 66 $0.6$ $90.0$ $55.7\%$ $78.6\%$ $63.7\%$ $47.8\%$ $68.9\%$ $84.2\%$
+$0.8$ $0.8$ 66 $0.6$ $180.0$ $50.2\%$ $74.0\%$ $55.2\%$ $41.5\%$ $63.7\%$ $83.1\%$
+A.4 Training & Inference Details
+We train our models using mixed precision to optimize both memory efficiency and training speed.
+The training process spans 100,000100{,}000 iterations, utilizing the Adam optimizer [27] with a OneCycle learning rate schedule [44], which gradually decays from a peak learning rate to enhance convergence.
+Consistent with RVT [18], we employ a mixed batching strategy to balance computational efficiency and memory usage. Specifically:
+  • Standard Backpropagation Through Time (BPTT): Applied to half of the training samples, allowing for full sequence training.
+  • Truncated BPTT (TBPTT): Used for the other half, reducing memory usage by splitting sequences into smaller segments.
+For data augmentation, we apply random horizontal flipping and zoom transformations (both zoom-in and zoom-out) to enhance the diversity of training samples.
+Our training process utilizes the YOLOX framework [63], a versatile object detection framework known for its efficient and high-performing architecture.
+We employ a multi-component loss function to optimize our model effectively:
+  • Intersection over Union (IoU) Loss: This loss component measures the overlap between the predicted bounding boxes and the ground-truth boxes, ensuring that the predicted regions closely match the actual object locations.
+  • Classification Loss: This component evaluates the accuracy of class predictions for each detected object, ensuring that the model correctly identifies the category of each detected instance.
+  • Regression Loss: This loss assesses the precision of the predicted bounding box coordinates, helping the model refine the location and size of bounding boxes to align closely with the ground-truth annotations.
+To ensure stable training, these loss components are averaged across both the batch and sequence length at each optimization step.
+This averaging process helps to reduce variance during training and facilitates smoother convergence of the model parameters.
+We also include a regularization term to balance the contribution of both modalities.
+Training Configuration.
+The training is conducted with a batch size of 8, which provides an optimal balance between efficient GPU utilization and memory requirements.
+Each training sample contains a sequence length of 11 frames, allowing the model to learn temporal dependencies effectively.
+The frame backbone’s weights are initialized using pre-trained ResNet, and the event backbone’s weights are initialized using pre-trained RVT.
+The learning rate is set to $1\times 10^{-4}$, following a OneCycle learning rate schedule that allows efficient exploration of the learning space and helps in achieving faster convergence.
+Hardware & Training Time.
+All training experiments are carried out on two NVIDIA RTX A50005000 GPUs, each with 2424GB of memory, providing the computational resources necessary for handling the high-resolution event data and RGB frames.
+The complete training process, including all iterations and model optimization, takes approximately one day, demonstrating the efficiency of our implementation in terms of both training speed and resource utilization.
+Appendix B Additional Quantitative Results
+B.1 Complete Results of Efficiency Analysis
+We include the complete results of the efficiency analysis in Tab. 8.
 B.2 Complete Results of Ablation Study
 We include the complete results of the ablation study in Tab. 9.
-B.3 Complete Results of Hyperparameter Searching
+B.3 Complete Results of Hyperparameter Search
 We include the complete results of the hyperparameter searching in Tab. 10.
-Appendix C C Additional Qualitative Results
-C.1 Visual Comparisons of Event Camera Detectors
-We include additional qualitative assessments in Fig. 7, Fig. 8, and Fig. 9.
-C.2 Visual Comparisons under Different Frequencies
-We include additional qualitative assessments in Fig. 10, Fig. 11, and Fig. 12.
-Appendix D D Potential Societal Impact & Limitations
+Appendix C Additional Qualitative Results
+C.1 Qualitative Comparisons of Event Detectors
+We present additional qualitative comparisons in Fig.
+9, Fig.
+10, and Fig.
+11 to further illustrate the advantages of FlexEvent over three state-of-the-art methods – RVT [18], SAST [39], and DAGr [16] – across nine diverse scenes.
+As shown in Fig.
+9, under fast-motion conditions, RVT and SAST fail to detect pedestrians, while DAGr misclassifies many distant objects as pedestrians or vehicles.
+In contrast, FlexEvent accurately captures all objects, demonstrating superior robustness in challenging high-speed scenarios.
+Similarly, in Fig.
+10, RVT frequently misses objects in cluttered scenes, and both SAST and DAGr mistakenly recognize pedestrians and distant vehicles under noisy, occluded conditions.
+Here again, FlexEvent excels by reliably detecting all relevant objects, highlighting its ability to handle complex urban environments.
+Finally, in Fig.
+11, rapidly changing illumination leads RVT and DAGr to misinterpret subtle motion cues; yet FlexEvent preserves stable performance, underscoring its enhanced temporal awareness.
+These observations confirm that FlexEvent not only surpasses competing methods in quantitative benchmarks but also consistently delivers high detection accuracy in real-world conditions – ensuring reliable performance in fast-moving, cluttered, or dynamically lit scenes.
+C.2 Comparisons under Different Frequencies
+We further evaluate the performance of FlexEvent across diverse event frequencies in Fig.
+12, Fig.
+13, and Fig.
+14.
+Specifically, we compare our method with RVT under settings ranging from low-frequency ($20$ Hz) to high-frequency ($180$ Hz) event streams, including an extreme scenario with empty events.
+In Fig.
+12, RVT struggles to detect distant or partially occluded objects at both low and high frequencies, whereas FlexEvent consistently identifies all targets by adaptively fusing frame-based evidence with event cues.
+Similarly, in Fig.
+13, RVT suffers substantial performance drops under sparse event conditions, often overlooking pedestrians; yet FlexEvent maintains reliable detection through its frequency-adaptive training strategy.
+Finally, in Fig.
+14, even when event input is minimal or missing, FlexEvent retains high accuracy by leveraging complementary frame information.
+These comparisons reaffirm our robustness and adaptability, highlighting our ability to handle a broad spectrum of event frequencies while preserving superior detection performance.
+Appendix D Potential Societal Impact & Limitations
 In this section, we discuss the potential societal impact of FlexEvent, including its positive contributions, broader implications, and known limitations.
@@
 For example, autonomous vehicles equipped with our approach can better detect pedestrians, cyclists, and other vehicles in real time, potentially reducing accidents and saving lives.
-Additionally, the computational efficiency provided by the adaptive event-frame fusion (FlexFuser) and frequency-adaptive learning (FAL) mechanisms reduces the need for resource-intensive training processes.
+Additionally, the computational efficiency provided by the adaptive event-frame fusion (FlexFuse) and frequency-adaptive learning (FlexTune) mechanisms reduces the need for resource-intensive training processes.
 This contributes to the broader societal goal of making advanced AI technologies more accessible and less energy-intensive, thereby minimizing the environmental impact of large-scale AI models.
@@
 D.3 Known Limitations
-While FlexEvent demonstrates significant performance improvements, there are several known limitations to our approach.
-Dependence on Event Camera Quality.
+While FlexEvent demonstrates significant performance improvements, there are several known limitations to our approach, which can be summarized as follows:
+Dependence on the Quality of Event Camera Data.
 The effectiveness of our approach relies on the quality of the event camera sensor.
@@
 Resource Requirements for High-Frequency Data.
-While FlexFuser mitigates the computational cost of training on high-frequency event data, processing extremely high-frequency event streams still requires substantial computational resources during inference.
+While FlexFuse mitigates the computational cost of training on high-frequency event data, processing extremely high-frequency event streams still requires substantial computational resources during inference.
 This could limit the scalability on resource-constrained devices or in real-time applications with stringent latency requirements.
+Appendix E Public Resources Used
+In this section, we acknowledge the public resources used, during the course of this work.
+E.1 Public Datasets Used
+  • DSEC [^2] https://dsec.ifi.uzh.ch........................................................................................................................................................................CC BY-SA 4.0
+  • DSEC-Det [^3] https://github.com/uzh-rpg/dsec-det........................................................................................................................................................................GNU General Public License v3.0
+  • DSEC-Detection [^4] https://github.com/abhishek1411/event-rgb-fusion........................................................................................................................................................................Creative Commons Zero v1.0 Universal
+  • DSEC-MOD [^5] https://github.com/ZZY-Zhou/RENet........................................................................................................................................................................Unknown
+  • Gen 1666https://www.prophesee.ai/2020/01/24/prophesee-gen1-automotive-detection-dataset........................................................................................................................................................................Prophesee Gen1 Automotive Detection Dataset License
+  • 1 Mpx [^7] https://www.prophesee.ai/2020/11/24/automotive-megapixel-event-based-dataset........................................................................................................................................................................Prophesee 1Mpx Automotive Detection Dataset License
+E.2 Public Implementations Used
+  • RVT [^8] https://github.com/uzh-rpg/RVT........................................................................................................................................................................MIT License
+  • SAST [^9] https://github.com/Peterande/SAST........................................................................................................................................................................MIT License
+  • SSM [^10] https://github.com/uzh-rpg/ssms_event_cameras........................................................................................................................................................................Unknown
+  • LEOD [^11] https://github.com/Wuziyi616/LEOD........................................................................................................................................................................MIT License
+  • DAGr [^12] https://github.com/uzh-rpg/dagr........................................................................................................................................................................GNU General Public License v3.0
+  • RENet [^13] https://github.com/ZZY-Zhou/RENet........................................................................................................................................................................Unknown
 Figure 9: Additional qualitative results of state-of-the-art event camera detectors.
-We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det.
+We compare the proposed FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det.
 Best viewed in colors.
 Figure 10: Additional qualitative results of state-of-the-art event camera detectors.
-We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det.
+We compare the proposed FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det.
 Best viewed in colors.
 Figure 11: Additional qualitative results of state-of-the-art event camera detectors.
-We compare the proposed FlexEvent with RVT (Gehrig & Scaramuzza, 2023), SAST (Peng et al., 2024), and DAGr (Gehrig & Scaramuzza, 2024) on the test set of DSEC-Det.
+We compare the proposed FlexEvent with RVT [18], SAST [39], and DAGr [16] on the test set of DSEC-Det.
 Best viewed in colors.
-Figure 10: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario.
+Figure 12: Additional qualitative comparisons of the RVT model [18] and the proposed FlexEvent under different event camera operation frequencies (2020 Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario.
 The experiments are conducted on the test set of DSEC-Det.
 Best viewed in colors.
-Figure 11: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario.
+Figure 13: Additional qualitative comparisons of the RVT model [18] and the proposed FlexEvent under different event camera operation frequencies (2020 Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario.
 The experiments are conducted on the test set of DSEC-Det.
 Best viewed in colors.
-Figure 12: Additional qualitative comparisons of the RVT model (Gehrig & Scaramuzza, 2023) and the proposed FlexEvent under different event camera operation frequencies ($20$ Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario.
+Figure 14: Additional qualitative comparisons of the RVT model [18] and the proposed FlexEvent under different event camera operation frequencies (2020 Hz, $60$ Hz, $90$ Hz, and $180$ Hz) and the empty event scenario.
 The experiments are conducted on the test set of DSEC-Det.
 Best viewed in colors.
-Appendix E E Public Resources Used
-In this section, we acknowledge the public resources used, during the course of this work.
-E.1 Public Datasets Used
-E.2 Public Implementations Used
+Acknowledgments
+The authors would like to sincerely thank the Program Chairs, Area Chairs, and Reviewers for the time and effort devoted during the review process.